### v2

# Model-Agnostic, Temperature-Informed Sampling Enhances Cross-Year Crop Mapping with Deep Learning

Abstract
Conventional benchmarks for crop type classification from optical satellite time series typically assume access to labeled data from the same year and rely on fixed calendar-day sampling. This limits generalization across seasons, where crop phenology shifts due to interannual climate variability, and precludes real-time application when current-year labels are unavailable. Furthermore, uncertainty quantification is often neglected, making such approaches unreliable for crop monitoring applications. Inspired by ecophysiological principles of plant growth, we propose a simple, model-agnostic sampling strategy that leverages growing degree days (GDD), based on daily average temperature, to replace calendar time with thermal time. By uniformly subsampling time series in this biologically meaningful domain, the method emphasizes phenologically active growth stages while reducing temporal redundancy and noise. We evaluate the method on a multi‐year Sentinel-2 dataset spanning all of Switzerland, training on one growing season and testing on other seasons. Compared to state‐of‐the‐art baselines, our method delivers substantial gains in classification accuracy and, critically, produces more calibrated uncertainty estimates. Notably, our method excels in low-data regimes and enables significantly more accurate early-season classification. With only 10 percent of the training data, our method surpasses the state-of-the-art baseline in both predictive accuracy and uncertainty estimation, and by the end of June, it achieves performance similar to a baseline trained on the full season. These results demonstrate that leveraging temperature data not only improves predictive performance across seasons but also enhances the robustness and trustworthiness of crop‐type mapping in real‐world applications.
keywords:
Crop Mapping, Deep Learning, Transformer, Temperature Data, Time Series, Uncertainty Quantification
1 Introduction
Monitoring agricultural land use is crucial for the sustainability of our food systems (Gómez et al.,, 2016). As the world’s population grows, climates shift, and diets evolve, there is a pressing need for monitoring tools that deliver reliable, affordable, and scalable insights into farming practices. In contrast to traditional methods, such as field surveys and self-reported data, which can be labor-intensive, expensive, and susceptible to errors, remote sensing-based approaches offer a faster, more cost-effective, and less subjective alternative for tracking land-use dynamics (Weiss et al.,, 2020).
Recent progress in remote sensing technologies and machine learning algorithms has significantly enhanced the ability to monitor agricultural land use. High-resolution, multi-spectral data from platforms such as Sentinel-2 provide frequent, detailed observations of croplands at minimal cost. By applying advanced machine learning methods, including conventional ensemble learners, e.g, XGBoost (Chen and Guestrin,, 2016) and deep neural networks, e.g, Transformers (Vaswani et al.,, 2017), to these time series images, we can capture crop-specific growth dynamics and distinguish among diverse management practices with high precision. These integrated approaches enable automated, scalable mapping of crop types and growth stages, supporting timely, data driven decisions for sustainable agriculture (Turkoglu,, 2023).
Figure 1: Cumulative Growing Degree Days (GDD) across Switzerland for multiple years.
Modern deep learning methods (Garnot and Landrieu, (2021), Turkoglu et al., 2021a , Metzger et al., (2021), Tarasiou et al., (2023), Rußwurm and Körner, (2018)) excel in classifying crop types from optical satellite image time series. However, these models are typically trained on curated datasets with limited size and/or validated on data collected within a single growing season (Rußwurm et al., (2019), Garnot and Landrieu, (2021), Schneider et al., (2023), Turkoglu et al., 2021a ), neglecting interannual shifts in growth driven by climate variability and thus limiting their applicability across years and regions. Furthermore, their dependence on current-year labeled data for training precludes real time deployment during the ongoing season. As also noted by Roscher et al., (2024), model development should focus on datasets with real-world relevance to further improve accuracy, generalization ability, and real impact on end user applications. Lastly, although these approaches may achieve state-of-the-art accuracy, they rarely incorporate uncertainty quantification, leaving predictive confidence unassessed and often uncalibrated. Ensuring the models are well-calibrated makes them reliable components of agricultural monitoring systems, where miscalibration could otherwise lead to costly resource misallocations or risk mismanagement.
Figure 2: NDVI time series of two parcels with the same crop rotation over two years (2021 and 2022), aligned by calendar date (left) and by cumulative growing degree days (GDD) (right).
Seen from an ecophysiological point of view, crop development is majorly driven by thermal conditions, since temperature plays a pivotal role in determining growth and phenological development. Temperature can vary substantially year-to-year within the same region or across different regions, as illustrated in Figure 1. It shows how the annual cumulative daily temperature in Switzerland varies between seasons. Similarly, Figure 2 (left) depicts Normalised Difference Vegetation Index (NDVI) time series for a single field over two years, revealing a noticeable shift in the crop’s growth trajectory despite the crop type remaining constant. Also see Figure 3, which illustrates how the same agricultural landscape can appear markedly different in satellite imagery across different years. By expressing crop development relative to a thermal time calendar, using growing degree days (GDD) rather than calendar days, the apparent interannual mismatch in growth cycles becomes significantly less pronounced (Figure 2 (right)). This motivates integrating temperature-based normalization into crop monitoring workflows to enhance model robustness and generalization across years. While other environmental drivers like precipitation undoubtedly influence crop growth, their incorporation falls beyond the scope of this study and is reserved for future research.
Figure 3: Year-to-year variation of the same agricultural landscape on June 15th from 2021 to 2023.
In this work, we build upon the state-of-the-art attention-based deep learning architecture proposed by Garnot and Landrieu, (2021), and introduce a simple, model-agnostic sampling strategy that incorporates daily average temperature to reduce redundancy and noise in satellite image time series. Deep learning models for crop classification typically rely on subsampled time series, as crop development progresses gradually and many satellite observations are temporally redundant. Subsampling is also necessary to reduce sequence length for efficient model training and inference within limited GPU memory. To better guide this selection, we replace calendar days with growing degree days (GDD) as the temporal reference and perform uniform subsampling in this thermal time domain. This approach emphasizes the crop’s biologically active periods, ensuring that the model focuses on the most informative observations.
We evaluate our GDD-informed sampling approach on a multi-year, country-wide dataset spanning three Swiss growing seasons by training on one year and testing on the other two. We benchmark against state-of-the-art baselines, evaluate low-data scenarios using just 10% of training samples, assess early-season classification performance, and examine uncertainty calibration for reliable decision support. We hypothesize that GDD-informed sampling will boost both classification accuracy and uncertainty estimation, also including under data scarcity and in early-growth phases, thereby addressing critical robustness and timeliness challenges in operational crop monitoring.
In summary, our contributions are as follows:
  • We introduce a simple yet effective model-agnostic approach that takes into account ecophysiological principles of plant growth: Thermal-Time-based Temporal Sampling (T3S), which allows integration of temperature data into any deep learning model, irrespective of model architecture.
  • We benchmark the T3S approach to other state-of-the-art approaches. The results demonstrate substantial improvements in predictive accuracy and uncertainty quantification across extensive country-wide, cross-year experiments, including low-data regimes and early-season classification settings.
  • We publicly release the SwissCrop dataset, a comprehensive, multi-year crop dataset, along with code to support further research and facilitate seamless integration into deep learning workflows.
2 Related Work
Deep Learning for Crop Mapping from Satellite Imagery.
Deep learning has dramatically improved crop‐type classification from satellite time series, outperforming traditional methods based on handcrafted features (Turkoglu,, 2023). Early sequence‐modeling approaches used recurrent architectures to capture phenological dynamics. Long short‐term memory (LSTM) networks demonstrated strong performance on multi‐temporal inputs (Rußwurm and Körner,, 2017), and the convolutional‐RNN (convRNN) version further boosted both accuracy and efficiency (Turkoglu et al., 2021a, ).
Attention mechanisms have emerged as a powerful alternative for modeling long‐range temporal dependencies. Rußwurm and Körner, (2020) and Garnot et al., (2020) were the first to apply self‐attention to crop classification from Sentinel-2 data and showed improved results. Series, (2020) then proposed the lightweight Temporal Attention Encoder (L-TAE), and Garnot and Landrieu, (2021) embedded L-TAE within a U-Net backbone to adaptively weigh temporal features across space and time. More recently, fully attention-based Vision Transformer (ViT) adaptations have been explored for crop mapping (Bai et al.,, 2023), and Qin et al., (2024) leveraged the recently proposed Mamba state‐space architecture to improve spatio‐temporal classification performance.
To incorporate domain knowledge in deep learning architectures, several works introduce structured priors and hierarchical labels. Turkoglu et al., 2024b exploited multi‐scale crop class hierarchies within a parameter‐efficient convRNN (i.e, convSTAR (Turkoglu et al., 2021b, )) framework, and prototype‐based supervision has been used to integrate class relationships into metric learning (Sainte Fare Garnot and Landrieu,, 2020). Positional encodings tailored to agricultural cycles have also been proposed: (Nyborg et al.,, 2022) introduced thermal-time encoding, using cumulative growing degree days, to normalize phenological shifts across regions. This is the most relevant research for this paper. Moreover, practical deployment demands handling early-season data and irregular revisit intervals. Rußwurm et al., (2023) added an early classification reward loss to LSTM training, improving accuracy in the season when only partial time series are available. Metzger et al., (2021) leveraged neural ordinary differential equations (NODE) to process noisy and irregularly sampled time series, and Cai et al., (2023)’s collect–update–distribute design further addresses sampling irregularity.
Multi-Year Crop Mapping from Satellite Imagery.
Multi-year crop mapping faces challenges due to temporal domain shifts such as interannual variability in weather, phenological differences, and sensor calibration issues. These factors often degrade the performance of models trained on single-season data when applied across multiple years. Turkoglu et al., 2024a extended the experiments of Turkoglu et al., 2021a to a multi‐year setting, training and testing on data from two different years, and reporting preliminary results. Blickensdörfer et al., (2022) employed a Random Forest classifier using dense Sentinel-2 and Landsat-8 time series, monthly Sentinel-1 composites, and environmental variables to evaluate the relative contributions of optical, radar, and environmental data across three meteorologically distinct years (2017–2019). Dubrovin et al., (2024) demonstrated that reconstructing multi-year Sentinel-2 NDVI time series with Fourier-based fitting and applying Gradient Boosting enables improved crop mapping across different years. Gao et al., (2023) investigated ways to optimize the use of multi-year samples in a within-season crop classification model. Qi et al., (2024) presented a Google Earth Engine workflow that fuses field-sample data with multi-year Sentinel-2 time series to map cropping systems on smallholder farms, achieving consistently robust results across diverse landscapes. Li et al., (2023) benchmarked deep learning segmentation models (e.g., U-Net and SegFormer (Xie et al.,, 2021)) for crop mapping across multiple years, while Sykas et al., (2022) evaluated convLSTM and convSTAR (Turkoglu et al., 2021b, ) networks in a similar multi-year context. Very recently, Khallaghi et al., (2025) shows improved results for cross-year crop mapping combining a U-Net with photometric augmentation, Tversky-Focal loss, and Monte Carlo (MC) Dropout.
Uncertainty Modeling in Deep Learning.
Predictive uncertainty is commonly split into aleatoric uncertainty, which stems from inherent data variability such as class overlap or sensor noise, and epistemic uncertainty, which reflects model weight uncertainty due to insufficient coverage of the input space (Kendall and Gal,, 2017). Exact Bayesian inference in deep networks is generally intractable, leading to approximate methods like variational inference (e.g., Bayes by Backprop,  Blundell et al., 2015), Markov Chain Monte Carlo techniques such as Stochastic Gradient Langevin Dynamics (Welling and Teh,, 2011). However, these approaches often struggle to scale to large architectures (Gustafsson et al.,, 2019). Probabilistic ensembles, particularly deep ensembles, currently represent the gold standard for epistemic uncertainty estimation in deep networks, providing well‐calibrated and robust predictions by aggregating outputs from multiple independently trained models (Lakshminarayanan et al.,, 2017). However, the need to train and store several large networks incurs substantial computational and memory overhead. To address this, implicit ensemble techniques, such as Monte Carlo (MC) Dropout (Gal and Ghahramani,, 2016), Snapshot Ensembles (Huang et al.,, 2017), Masksembles (Durasov et al.,, 2021), and FiLM-Ensemble (Turkoglu et al.,, 2022), approximate ensemble diversity via stochastic weight perturbations or dynamic weight combinations, significantly reducing resource requirements. Despite these gains in efficiency, implicit methods generally underperform full ensembles in both predictive accuracy and uncertainty calibration. More recently, LoRA-Ensemble has been introduced as a parameter‐efficient alternative for transformer‐based architectures, achieving uncertainty estimates competitive with deep ensembles at a fraction of the cost (Mühlematter et al.,, 2024).
Uncertainty Modeling in Satellite Imagery and Crop Mapping
Uncertainty quantification has become a standard component of satellite-based land-cover and crop-mapping workflows, as it provides essential information for risk-aware decision-making in agricultural management and ecosystem monitoring. Generating per-pixel or per-field uncertainty maps alongside deterministic predictions enables practitioners to identify low-confidence areas and tailor downstream analyses accordingly. For example, Lang et al., (2022) employs a deep ensemble of convolutional neural networks to estimate global canopy height from sparse LiDAR samples and multi-sensor satellite imagery, producing uncertainty intervals that reflect both model and data variability. Becker et al., (2023) extends this ensemble approach to country-scale forest structure estimation at 10m resolution, demonstrating that dense uncertainty fields help flag regions where auxiliary field data are most needed.
In the context of crop yield and type mapping, several recent studies have adopted stochastic inference techniques. Tian et al., (2024) introduces an attention-guided multi-level crop network for winter wheat yield estimation at the county scale, applying MC Dropout during inference to derive spatially explicit uncertainty maps of yield forecasts. Similarly, Khallaghi et al., (2025) leverages MC Dropout in a U-Net segmentation framework to quantify per-pixel uncertainty in annual cropland masks across multiple years in Ghana, using the resulting confidence estimates to discard unreliable predictions. Liu et al., (2024) integrates MC Dropout in a ViT–ResNet feature encoder to produce field-level uncertainty scores, filtering out low-confidence labels to improve overall mapping precision. Beyond dropout-based methods, Baumert et al., (2024) proposes a probabilistic crop-type mapping method in which an ensemble of classifiers, each trained with perturbed inputs and hyperparameters on Sentinel-1, Sentinel-2, and Landsat-8 data, generates multiple crop maps over Germany; per-pixel frequency counts across the ensemble yield robust probability distributions and uncertainty metrics. Maleki et al., (2024) refines the 30m cropland data by harnessing its built-in Random Forest confidence layer: low-certainty regions are iteratively reclassified using targeted ancillary datasets, resulting in a more accurate and self-aware mapping product. Collectively, these approaches demonstrate the critical role of uncertainty modeling in improving the reliability and interpretability of satellite-driven crop and canopy mapping systems.
3 Method
Our method builds on the U-Net with Temporal Attention Encoder (U-TAE) (Ronneberger et al.,, 2015; Garnot and Landrieu,, 2021), a state-of-the-art deep learning model for dense segmentation of satellite image time series. Prior to presenting our sampling approach, we examine U-TAE on a standardized benchmark by adjusting sequence lengths and positional encodings to identify the minimal temporal resolution and ordering information needed for high classification accuracy. These insights directly inform our design choices and motivate the development of Thermal-Time-based Temporal Sampling (T3S), a model-agnostic method that aligns input selection with crop physiology.
Figure 4: Learning curves of the U-TAE model with different position encodings (PE) on the PASTIS dataset, showing training loss, training accuracy, validation loss, and validation accuracy respectively.
3.1 U-Net with Temporal Attention Encoder (U-TAE)
The U-TAE model is a spatio-temporal neural network designed to extract dense and temporally-aware features from satellite image time series. It consists of a shared 2D convolutional encoder applied to each timestamp independently, followed by a temporal self-attention module that aggregates the encoded features across time. This attention module uses positional encoding to preserve temporal ordering and applies scaled dot-product attention to dynamically weigh each timestamp based on its relevance. The aggregated multi-temporal representation is then decoded through a U-Net-style decoder with skip connections, enabling high-resolution semantic segmentation. The architecture can handle variable-length input sequences and produce temporally adaptive representations of land cover. Refer to Garnot and Landrieu, (2021) for more details.
3.2 Preliminary Analysis with U-TAE
Subsampling the satellite image time series is necessary for two main reasons:
  • Redundancy and noise in temporal observations. There is currently no comprehensive study establishing the exact number of time steps required to classify crops reliably using U-TAE or other architectures. Existing works vary widely, for example, Garnot and Landrieu, (2021) uses up to 61 observations, while Turkoglu et al., 2021a employs 71 timestamps. In principle, Sentinel-2 can provide an observation every 2 to 3 days over Switzerland, making very long sequences feasible. However, such long sequences are often redundant and noisy: Metzger et al., (2021) reports that roughly 50% of observations can be unusable due to cloud cover. Moreover, crops may develop relatively slowly, e.g. in periods with low temperatures. Consequently, analogous to a slowly evolving video, densely sampling in time adds little new information for distinguishing crop types.
  • Computational efficiency and memory constraints. Deep temporal models like U-TAE require substantial GPU memory when processing long sequences. To maintain a reasonable batch size during training, it is essential to limit the number of timestamps. Note that larger batches reduce the variance in the estimation of the gradient, resulting in a stronger and more stable training signal (Räisä et al.,, 2024). This is particularly advantageous for highly imbalanced datasets: choosing a batch size several times the minority-to-majority class ratio ensures that each minibatch contains sufficient minority-class examples to learn meaningful features (Google Developers,, 2025).
To ensure our evaluation remains unbiased, we base our initial analysis on the well-established PASTIS benchmark, a concise and reproducible toy crop classification dataset introduced alongside U-TAE by Garnot et al. (2021). By mirroring SwissCrop’s multi-temporal structure while remaining fully independent, PASTIS enables rapid iteration and guarantees that insights from our analysis directly inform algorithm design without the dataset itself biasing the method.
In this dataset each time series contains on average 48 Sentinel-2 observations (up to 61). We uniformly subsampled each sequence to 24 time steps, and trained with a batch size of 4 (see Garnot and Landrieu, (2021) for full training details). As shown in Table 1, this simple uniform subsampling yields only a 0.2% absolute drop in overall accuracy, despite no selective filtering of cloudy or low-information dates. Crucially, GPU memory usage is cut in half, and the number of multiply–accumulate operations (MACs) per forward pass with a single batch is reduced around 50%, dramatically lowering computational cost and freeing capacity for larger batches or faster training and inference. Accordingly, we set the temporal length to 24 for the experiments.
Table 1: Performance comparison for varying temporal length on the Pastis dataset. Results are averaged over a 5-fold experimental setting of the Pastis dataset. Refer to Garnot and Landrieu, (2021) for details. M stands for GPU memory.
$T_{min}$ $T_{max}$ $T_{mean}$ M(Gb) MACs(B) Acc(%)
38 61 48 16.8 288 $83.1\pm 0.5$
24 24 24 8.0 145 $82.9\pm 0.5$
Positional encoding (PE) injects information about each observation’s position in the sequence into the attention mechanism, enabling the network to distinguish the order and spacing of time steps. In U-TAE, we add a standard PE, computed as a set of sinusoids at different frequencies, to the input embeddings to mark each time index. We experimented with alternative encodings that explicitly encode absolute time intervals rather than simple positional indices, but found no performance gain, indicating that U-TAE primarily leverages the relative ordering of observations rather than their exact timestamps (Figure 4).
Specifically, we trained three U-TAE variants with different position encodings:
  • noPE: All position embeddings are set to zero.
  • linearPE: Positions are encoded as sequential integers (1, 2, 3, …).
  • calendarPE: Positions are encoded as the actual calendar dates of each observation (1, …, 365).
As shown in Figure 4, the noPE model underperforms in both training and validation metrics, confirming that some ordering information is necessary. However, the linearPE and calendarPE models converge to nearly identical performance, suggesting that incorporating absolute time values (whether calendar dates or temperature‐derived values) does not yield additional benefits beyond preserving the relative order of observations. This result led us to conclude that using thermal time directly as a positional encoding is unlikely to be very effective, as we show in our experiments.
In summary, our preliminary findings reveal that (i) substantial temporal subsampling incurs only a marginal accuracy loss while reducing both GPU memory usage and computational cost drastically, and (ii) U-TAE relies chiefly on the relative ordering of observations, whereas absolute timestamps offer limited benefit. These insights suggest that, rather than encoding raw dates, we should focus on selecting the most informative, developmentally relevant observations. In the next section, we leverage this principle to introduce model‐agnostic approach that partitions input sequences according to growing degree days to better align sampling with crop physiology.
3.3 T3S: Thermal-Time-based Temporal Sampling
Satellite-based crop monitoring approaches often rely on fixed-interval calendar-day sampling to construct input sequences. However, this overlooks the ecophysiological processes of plant growth such as their dependence on temperature. As crop development is tightly coupled with temperature, we propose a simple yet effective model-agnostic sampling method, Thermal-Time-based Temporal Sampling (T3S), which replaces fixed calendar intervals with intervals defined in thermal time, using Growing Degree Days (GDD) as the temporal reference.
GDD provides a cumulative measure of heat accumulation, reflecting the physiological development of crops more faithfully than absolute time. For a given day $d$, GDD is computed as:
$\mathrm{GDD}_{d}=\max\left(0,\frac{T_{\text{max},d}+T_{\text{min},d}}{2}-T_{\text{base}}\right),$ (1)
where $T_{\text{max},d}$ and $T_{\text{min},d}$ are the daily maximum and minimum temperatures, respectively, and $T_{\text{base}}$ is a base threshold below which crop development is considered negligible. In our experiments, we use $T_{\text{base}}=0^{\circ}$, a commonly used value for temperate climate crops.
We then compute the cumulative GDD over time:
$\mathrm{GDD}_{D}=\sum_{d=1}^{D}\mathrm{GDD}_{d}.$ (2)
To construct a time series of $L$ observations, we partition the full GDD range $\left[\mathrm{GDD}{\min},\mathrm{GDD}{\max}\right]$ into $L$ uniform intervals. Within each interval, we select the least cloudy satellite observation, prioritizing data quality. Importantly, Sentinel-2 products include a freely available cloud mask layer, allowing direct use of cloud information without additional preprocessing. This ensures that the selected timestamps correspond to ecophysiological meaningful intervals that are consistent across different years, even when calendar dates diverge significantly due to seasonal variability.
The T3S algorithm is outlined in Algorithm 1. It takes as input the full set of timestamps, the cumulative thermal time values, i.e, GDD, the satellite data cube, a cloud mask for each observation, and the desired number of samples $L$. For each thermal interval, it identifies candidate observations and selects the one with the fewest cloud-contaminated pixels. The final set of selected indices is sorted to preserve temporal order. Figure 5 illustrates the overall workflow of our approach.
This sampling method offers several advantages. First, it aligns input sequences more closely with crop development, reducing temporal noise caused by inter-annual climatic shifts. Second, it improves data efficiency by avoiding redundant or uninformative observations. Lastly, T3S is agnostic to model architecture and can be seamlessly integrated into any time series-based crop monitoring pipeline.
Figure 5: The first row shows conventional uniform sampling; the second row shows uniform sampling where the least cloudy observation is chosen in each interval (referred to as “Deformable” baseline); the third row shows the proposed Thermal‐Time‐based Temporal Sampling (T3S) approach. In T3S, each square box represents the temperature sum accumulated since the previous one, illustrating intervals with equal thermal time.
Algorithm 1 T3S
1:function T3S(timeStamps, GDD, data, cloudMask, $L$)
2:     $t_{\min}\leftarrow\min(\text{GDD})$
3:     $t_{\max}\leftarrow\max(\text{GDD})$
4:     $\Delta\leftarrow(\,t_{\max}-t_{\min}\,)/L$   #$L$ is target temporal length
5:     $selected\leftarrow[\ ]$
6:     $current\leftarrow t_{\min}$
7:     while $|selected|<L$ and $current<t_{\max}$ do
8:         $next\leftarrow current+\Delta$
9:         $window\leftarrow\{\,i\mid current\leq\text{GDD}[i]<next\}$
10:         if $window\neq\emptyset$ then
11:              $best\leftarrow\displaystyle\arg\max_{i\in window}\sum(\text{cloudMask}[i]=0)$
12:              append $best$ to $selected$
13:         end if
14:         $current\leftarrow next$
15:     end while
16:     sort($selected$)
17:     return data[:, selected, …]
18:end function
3.4 Training Details
The model was trained for 100 epochs using a batch size of 8 and a fixed temporal length of 24. We employed the Adam optimizer with an initial learning rate of 0.001, scheduled by a OneCycleLR policy with a maximum learning rate of 0.01 and a warm-up phase covering 30% of the total training steps. The U-TAE architecture is configured with an encoder having channel widths $[128,128,128,256]$, a symmetric decoder with widths $[32,64,128,256]$ and strided convolutions with kernel size $k=4$, stride $s=2$, and padding $p=1$. The temporal attention module employs $n_{\text{head}}=16$ heads and a model dimension of $d_{\text{model}}=256$. Input features are standardized to zero mean and unit variance on a per-channel basis. We implemented the model in PyTorch and trained it on a single NVIDIA RTX 4090 GPU over roughly three days. The full codebase and pretrained checkpoints will be publicly available on GitHub.
4 Dataset
In this work, we work with SwissCrop 2021, 2022, 2023, a three-year crop dataset covering the entirety of Switzerland from 2021 to 2023. SwissCrop combines Sentinel-2 Level-2A bottom-of-atmosphere multi-spectral imagery with annual crop‐type labels for each field. The dataset includes 50 distinct crop types, each representing the primary crop grown in a field during a given season. These labels where provided by farmers to the cantons of Switzerland according to a data model provided by the Federal Office for Agriculture  [^1] https://www.blw.admin.ch/de/landwirtschaftliche-kulturflaechen and joined together by the Swiss Centre for Agricultural Research, Agroscope. Unlike many existing datasets, e.g. Pastis dataset (Garnot and Landrieu,, 2021), SwissCrop faithfully reflects the pronounced class imbalance found in real‐world farming: a long‐tail distribution in which a handful of major crops dominate while the majority occur only sporadically (cf. Figure 6).
[…]
To complement these remote sensing observations, we incorporate high-resolution gridded climate data from MeteoSwiss, which supplies near-surface temperature fields (minimum, maximum, and average) on 1 km, 2 km, and 5 km grids (Federal Office of Meteorology and Climatology, MeteoSwiss,, 2024). In our study, we use the 1 km daily minimum (TminD) and maximum (TmaxD) temperature series to compute growing degree days (GDD; refer to Section 3.3).
We make the full SwissCrop dataset publicly available to the research community. We hope this resource will facilitate further advances in large-scale, cross-year crop classification, phenology modeling, and earth observation research.
Figure 6: SwissCrop dataset label distribution for the year 2021. Note the logarithmic scale of the y-axis.
Figure 7: Region of interest (top left), the SwissCrop dataset (bottom left) and its coverage by data cubes (center), with an example of a single analysis-ready cube (right).
5 Experiment & Results
We evaluate the proposed T3S against four very relevant state-of-the-art baselines using a leave‐one‐year‐out scheme over the 2021–2023 SwissCrop dataset. In each of six cross‐year folds, two seasons are held out for testing while the remaining one season is used for training. To further assess robustness, we conduct a low‐data experiment using only 10% of the labels, and an early‐season study by truncating each test timeseries at the 50th and 75th time percentile (end of June and end of September respectively). We assess model predictive performance using Overall accuracy, the fraction of pixels correctly labeled; Intersection‐over‐Union (IoU), the ratio of the area of overlap between predicted and ground‐truth pixels to the area of their union; and Mean Intersection‐over‐Union (mIoU), averaged across classes. Calibration is measured by the Expected Calibration Error (ECE), which measures how far predicted confidences deviate from observed error rates, i.e., perfect calibration occurs when the estimated uncertainties match the actual likelihood of a misclassification. We also report Negative Log‐Likelihood (NLL), the mean negative log‐probability of the true class, and the Brier Score, the mean squared error between predicted probabilities and one‐shot labels.
Table 2: Performance comparison across six evaluation settings. The first two columns show the year of the training and test datasets. Best score for each setting and each metric in bold, second best underlined.
Train Test Method Accuracy % ($\uparrow$) mIoU % ($\uparrow$) IoU % ($\uparrow$) ECE ($\downarrow$) NLL ($\downarrow$) Brier Score ($\downarrow$)
2021 2022 U-TAE (Garnot and Landrieu,, 2021) 70.1 17.2 53.9 0.053 0.943 0.412
+MC-Dropout (Gal and Ghahramani,, 2016) 71.3 17.4 55.4 0.032 0.911 0.402
+Thermal-PE (Nyborg et al.,, 2022) 72.5 18.6 56.9 0.045 0.854 0.385
+Deformable Sampling 73.8 20.5 58.5 0.044 0.836 0.371
T3S (ours) 76.0 21.6 61.3 0.004 0.720 0.329
2021 2023 U-TAE (Garnot and Landrieu,, 2021) 71.4 16.7 55.6 0.052 0.947 0.399
+MC-Dropout (Gal and Ghahramani,, 2016) 72.5 17.0 56.8 0.037 0.918 0.394
+Thermal-PE (Nyborg et al.,, 2022) 71.3 15.5 55.4 0.035 0.925 0.401
+Deformable Sampling 74.6 18.9 59.6 0.032 0.822 0.359
T3S (ours) 77.6 21.0 63.5 0.005 0.691 0.307
2022 2021 U-TAE (Garnot and Landrieu,, 2021) 70.8 16.2 54.8 0.035 0.936 0.408
+MC-Dropout (Gal and Ghahramani,, 2016) 71.2 16.2 55.3 0.024 0.913 0.399
+Thermal-PE (Nyborg et al.,, 2022) 72.8 18.2 57.2 0.077 1.022 0.403
+Deformable Sampling 76.8 20.4 62.3 0.043 0.833 0.339
T3S (ours) 77.3 20.1 63.0 0.013 0.697 0.310
2022 2023 U-TAE (Garnot and Landrieu,, 2021) 72.1 17.4 56.4 0.041 0.875 0.389
+MC-Dropout (Gal and Ghahramani,, 2016) 72.8 17.2 57.2 0.028 0.853 0.381
+Thermal-PE (Nyborg et al.,, 2022) 63.9 14.2 46.9 0.123 1.374 0.525
+Deformable Sampling 76.0 20.8 61.4 0.073 0.936 0.363
T3S (ours) 76.8 20.9 62.4 0.019 0.708 0.318
2023 2021 U-TAE (Garnot and Landrieu,, 2021) 71.6 17.6 55.8 0.066 0.913 0.398
+MC-Dropout (Gal and Ghahramani,, 2016) 71.6 17.6 55.7 0.040 0.889 0.391
+Thermal-PE (Nyborg et al.,, 2022) 71.8 16.3 56.0 0.098 0.987 0.413
+Deformable Sampling 75.3 20.5 60.4 0.105 0.951 0.379
T3S (ours) 77.1 23.2 62.8 0.019 0.686 0.307
2023 2022 U-TAE (Garnot and Landrieu,, 2021) 71.2 18.4 55.3 0.059 0.957 0.410
+MC-Dropout (Gal and Ghahramani,, 2016) 72.0 18.6 56.2 0.043 0.929 0.400
+Thermal-PE (Nyborg et al.,, 2022) 75.3 19.6 60.4 0.085 0.929 0.366
+Deformable Sampling 75.8 21.9 61.0 0.081 0.827 0.355
T3S (ours) 76.9 22.3 62.5 0.006 0.687 0.311
Average U-TAE (Garnot and Landrieu,, 2021) $71.2\pm 0.6$ $17.3\pm 0.7$ $55.3\pm 0.8$ $0.051\pm 0.012$ $0.929\pm 0.028$ $0.403\pm 0.008$
+MC-Dropout (Gal and Ghahramani,, 2016) $71.9\pm 0.6$ $17.3\pm 0.7$ $56.1\pm 0.7$ $\underline{0.034}\pm 0.007$ $0.902\pm 0.029$ $0.395\pm 0.007$
+Thermal-PE (Nyborg et al.,, 2022) $71.3\pm 3.5$ $17.1\pm 1.9$ $55.5\pm 4.1$ $0.077\pm 0.029$ $1.015\pm 0.156$ $0.416\pm 0.053$
+Deformable Sampling $\underline{75.4}\pm 1.0$ $\underline{20.5}\pm 0.9$ $\underline{60.5}\pm 1.2$ $0.063\pm 0.025$ $\underline{0.868}\pm 0.045$ $\underline{0.361}\pm 0.014$
T3S (ours) $\textbf{77.0}\pm 0.5$ $\textbf{21.5}\pm 1.0$ $\textbf{62.6}\pm 0.7$ $\textbf{0.011}\pm 0.006$ $\textbf{0.698}\pm 0.012$ $\textbf{0.314}\pm 0.008$
Figure 8: Class-wise accuracy difference between the proposed T3S and U-TAE. Averaged over six experimental folds. Blue bars indicate classes where T3S outperforms U-TAE, while red bars indicate classes where T3S underperforms.
Baselines.
U-TAE is a U‐Net combined with a temporal attention encoder (i.e., self-attention layer) that ingests images at fixed calendar dates and uses sinusoidal day‐of‐year positional encodings; it serves as our primary reference. We combine U-TAE with MC Dropout (Gal and Ghahramani,, 2016): U-TAE+MC‐Dropout retains the same sampling grid but activates dropout layers at test time to perform Monte Carlo sampling, thereby approximating a Bayesian posterior and yielding pixel‐level uncertainty estimates without changing the input dates. Note that Monte Carlo Dropout remains one of the most popular and theoretically grounded methods for uncertainty calibration in deep learning (Kendall and Gal,, 2017). Empirically, the number of ensemble members is set to 5, and the drop rate is set to 0.2. U-TAE+Thermal-PE follows Nyborg et al., (2022) and replaces the sinusoidal calendrical encodings with absolute thermal‐time (Growing Degree Days) embeddings, giving the network explicit phenological context while still sampling on calendar dates. U-TAE+Deformable Sampling follows the same fixed‐interval calendar schedule (e.g. one image per fifteen days) but “deforms” each interval by selecting the least‐cloudy observation, thus reducing noise from clouds without altering the temporal grid.
Cross‐Year Generalization and Uncertainty Calibration.
T3S consistently outperforms all baselines in cross‐year generalization (see Table 2), delivering an approximate 6% absolute improvement in overall accuracy over U-TAE, alongside gains in IoU and mean IoU. The discrepancy between IoU and mean IoU reflects the dataset’s long‐tailed label distribution: rare classes contribute near-zero scores, pulling down the mean. Figure 8 shows that most crop categories benefit from T3S. Only the rarest classes (e.g., hop), for which both models lack sufficient training samples, remain at zero accuracy. Meadow accuracy drops slightly, reflecting U-TAE’s tendency to overpredict the dominant class and incur more false positives.
Crucially, T3S also yields much better‐calibrated outputs. T3S reduces ECE by roughly fourfold compared to U-TAE, approaching perfect calibration, and substantially lowers both Negative Log‐Likelihood and Brier Score, indicating sharper and more reliable uncertainty estimates. By comparison, U-TAE+MC‐Dropout achieves the second‐best calibration, modestly boosting accuracy and reducing ECE relative to vanilla U-TAE but still falling short of T3S. Note that MC-Dropout is specifically designed to improve the uncertainty calibration. U-TAE+Thermal-PE introduces phenological context via GDD embeddings and yields moderate gains in both performance and calibration, yet does not match the robustness of thermal‐time sampling. Refer to Section 3.2 for a detailed analysis of its limited efficacy. Deformable Sampling improves predictive accuracy through cloud filtering but does so at the expense of uncertainty quality, exhibiting higher ECE. Overall, these results highlight that aligning observations by thermal time via temporal sampling, combined with cloud‐aware selection yields the most robust and trustworthy cross‐year crop mapping models.
Qualitative Comparison.
Figure 11 contrasts T3S and U-TAE on both classification and uncertainty estimation, where pixel-wise uncertainty is defined as $1-\max(\mathrm{softmax})\in[0,1]$. In the first example, both models misclassify the highlighted region, but U-TAE remains overconfident in its incorrect label while T3S appropriately elevates uncertainty around the error. In the second example, T3S confines its errors to the fuzzy edges of the highlighted field, which is understandable given that farmers hand-draw plot boundaries and 10m pixels often cover two crop types, and its uncertainty map shows a pronounced hotspot exactly along those misclassified borders. In contrast, U-TAE not only mislabels the entire field but also produces a diffuse, misaligned uncertainty map that neither highlights its own errors nor respects the actual field geometry. In the third example, T3S again flags its error with a pronounced uncertainty hotspot, even revealing a likely ground-truth annotation mistake, while U-TAE provides no meaningful uncertainty signal. These examples illustrate that T3S not only reduces classification errors but also delivers well-calibrated uncertainty estimates that expose both model and dataset inconsistencies.
Low Data Regime.
In many agricultural settings, annotated crop labels are scarce or expensive to obtain, which makes low data performance a critical requirement for operational viability. To assess robustness under label scarcity, we conducted a six-fold experiment using only 10 % of the training data, randomly sampled once and held constant across methods, and evaluated both T3S and U-TAE on the full test set. As shown in Figure 9, both models incur drops in accuracy and increases in Expected Calibration Error when labels are limited, but U-TAE’s performance deteriorates far more rapidly. Remarkably, T3S trained on just 10 % of the data still outperforms U-TAE trained on the full dataset, and its calibration remains substantially more reliable.
This advantage stems from the incorporation of a meaningful inductive bias, namely the thermal time prior, which encodes domain knowledge about crop development into the model architecture. By constraining the hypothesis space to reflect biologically plausible growth patterns, T3S requires fewer examples to learn robust representations, making it far more data efficient and noise resilient, for example, against cloud cover and sensor artifacts, than models trained purely in a data-driven fashion. These results demonstrate that thermal time inductive bias confers strong resilience to small data scenarios, underscoring T3S’s practicality for real-world applications where labels are limited. See Section 6 for more discussion.
Early‐Season Classification.
Early season classification is an important aspect of this technology, since some applications require crop labeling during the growing season rather than after harvest. To evaluate T3S in this context, we truncate the test set time series at two percentiles, 75 percent (end of September) and 50 percent (end of June), and perform a six fold experiment in which both T3S and U-TAE are trained on full-season data and tested on the truncated series. Figure 10 shows that both models experience a decline in in-season performance, measured by predictive accuracy and uncertainty estimation, when using partial time series. At the 75 percent cutoff, U-TAE’s performance drops sharply, whereas T3S remains virtually unchanged. Moreover, at the 50 percent cutoff, T3S attains a similar level of accuracy and uncertainty calibration that U-TAE achieves using the full season of data. These results indicate that T3S not only improves in-season classification accuracy and uncertainty quantification but also underscores its practical suitability for operational in-season crop monitoring and decision-support applications, where timely and reliable predictions are critical.
Figure 9: Low data regime performance showing classification (a) accuracy and (b) Expected Calibration Error averaged over six folds when training on 10% of the data, comparing T3S (Ours) and U-TAE.
Figure 10: In‐season classification performance at the 75th‐ and 50th‐percentile cutoffs, corresponding to the end of September and the end of June, respectively, showing (a) accuracy and (b) Expected Calibration Error, averaged over six folds, for T3S (Ours) versus U-TAE.
Figure 11: Qualitative comparison for the proposed T3S and U-TAE. In performance maps, red indicates a wrong prediction while green represents a correct prediction. High uncertainty is indicated by yellow, while low is indicated with dark blue.
6 Discussion
Our experiments demonstrate that embedding GDD-informed sampling into a deep learning pipeline not only improves crop classification accuracy but also yields better-calibrated uncertainty estimates compared to leading baselines. This advantage grows when training data are scarce, highlighting the value of domain expertise under low-data regimes.
Injecting domain knowledge into a neural network, i.e., inductive bias, constrains the hypothesis space to solutions that ’make sense’ under known physical or biological rules and this prevents the model from assigning high confidence to predictions when evidence is weak or contradictory (Mitchell, (1980), Ahmadi Daryakenari et al., (2024)). In our case inductive bias is introduced via GDD-informed sampling. By incorporating thermal time alignment, we align satellite time series with actual crop phenology. This inductive bias acts as a regularizing prior, steering the model toward meaningful growth stages and away from noisy or cloud-contaminated observations. As a result, high-confidence predictions coincide with clear phenological cues, while lower-confidence outputs correctly signal uncertainty, demonstrating reliable probability calibration.
By combining deep neural networks with agronomic expertise, our approach not only enhances classification accuracy but also empowers the model to recognize and flag its own uncertainty, an essential feature in agricultural decision‐making, where misplaced confidence can have costly consequences.
7 Conclusion
In this paper, we introduced Thermal Time‐based Temporal
Sampling (T3S), a simple, model-agnostic method that uses cumulative growing degree days to select relevant observations within satellite time series for crop growth-related applications. Based on basic ecophysiological principles of crop growth, it reduces the computational load by removing redundant information and allows to prioritize relevant and undisturbed imagery. Although we evaluated T3S with attention-based deep learning architecture, it can be seamlessly integrated with any machine learning model. When compared to other state-of-the-art approaches, T3S not only enhances predictive accuracy but also markedly improves uncertainty estimation, an aspect often overlooked in the literature. Furthermore, we made evaluation on SwissCrop, a novel, large-scale, multi-year dataset spanning three growing seasons (2021–2023) across Switzerland’s agricultural landscapes, thereby providing a real-world benchmark for compared methods. By releasing both SwissCrop and our T3S implementation to the public, we establish a robust, reproducible benchmark for the community, facilitating fair comparisons and driving further innovation in crop‐growth modelling. For future work, incorporating additional environmental drivers such as precipitation could further strengthen temporal alignment and robustness across diverse agro-climatic regions.
8 Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
9 Acknowledgments
Mehmet Ozgur Turkoglu acknowledges support from Agroscope through research funding for the CropmAIpper project. Sélène Ledain acknowledges funding from the Swiss Federal Office of Agriculture and Agroscope within the Monitoring of the Swiss Agri-Environmental System (MAUS) program. This project was conducted in the context of the Swiss Agricultural Landscape Intelligence platform, funded through an Agroscope Research grant awarded to Helge Aasen.

### v3

@@
 Abstract
-Conventional benchmarks for crop type classification from optical satellite time series typically assume access to labeled data from the same year and rely on fixed calendar-day sampling.
-This limits generalization across seasons, where crop phenology shifts due to interannual climate variability, and precludes real-time application when current-year labels are unavailable.
-Furthermore, uncertainty quantification is often neglected, making such approaches unreliable for crop monitoring applications.
-Inspired by ecophysiological principles of plant growth, we propose a simple, model-agnostic sampling strategy that leverages growing degree days (GDD), based on daily average temperature, to replace calendar time with thermal time.
-By uniformly subsampling time series in this biologically meaningful domain, the method emphasizes phenologically active growth stages while reducing temporal redundancy and noise.
-We evaluate the method on a multi‐year Sentinel-2 dataset spanning all of Switzerland, training on one growing season and testing on other seasons.
-Compared to state‐of‐the‐art baselines, our method delivers substantial gains in classification accuracy and, critically, produces more calibrated uncertainty estimates.
-Notably, our method excels in low-data regimes and enables significantly more accurate early-season classification.
-With only 10 percent of the training data, our method surpasses the state-of-the-art baseline in both predictive accuracy and uncertainty estimation, and by the end of June, it achieves performance similar to a baseline trained on the full season.
-These results demonstrate that leveraging temperature data not only improves predictive performance across seasons but also enhances the robustness and trustworthiness of crop‐type mapping in real‐world applications.
-keywords:
-Crop Mapping, Deep Learning, Transformer, Temperature Data, Time Series, Uncertainty Quantification
+Crop type classification using optical satellite time series remains limited in its ability to generalize across seasons, particularly when crop phenology shifts due to inter-annual weather variability.
+This hampers real-world applicability in scenarios where current-year labels are unavailable.
+In addition, uncertainty quantification is often overlooked, which reduces the reliability of such approaches for operational crop monitoring.
+Inspired by ecophysiological principles of plant growth, we propose a simple, model-agnostic Thermal-Time-based Temporal Sampling (T3S) method that replaces calendar time with thermal time.
+By subsampling time series in this biologically meaningful way, our method highlights key periods within the growing season while reducing temporal redundancy and noise.
+We evaluate the T3S on a multi-year Sentinel-2 dataset covering the entirety of Switzerland, which allows us to assess all applied methods on unseen years.
+Compared to state-of-the-art baselines, our approach yields substantial improvements in classification accuracy and, critically, provides well-calibrated uncertainty estimates.
+Moreover, the T3S method excels in low-data regimes and enables significantly more accurate early-season classification.
+With just 10% of the training labels, it outperforms the current baseline in both accuracy and uncertainty calibration, and by the end of June, it achieves a performance similar to the full-season baseline model.
+The dataset and the code will be publicly available.
 1 Introduction
-Monitoring agricultural land use is crucial for the sustainability of our food systems (Gómez et al.,, 2016).
-As the world’s population grows, climates shift, and diets evolve, there is a pressing need for monitoring tools that deliver reliable, affordable, and scalable insights into farming practices.
-In contrast to traditional methods, such as field surveys and self-reported data, which can be labor-intensive, expensive, and susceptible to errors, remote sensing-based approaches offer a faster, more cost-effective, and less subjective alternative for tracking land-use dynamics (Weiss et al.,, 2020).
+Monitoring agricultural land use is crucial for the sustainability of our food systems [16].
+As the world’s population grows, climates becomes extremer, and diets evolve, there is a pressing need for monitoring tools that deliver reliable, affordable, and scalable insights into farming practices.
+In contrast to traditional methods, such as field surveys and self-reported data, which can be labor-intensive, expensive, and susceptible to errors, remote sensing-based approaches offer a faster, more cost-effective, and less subjective alternative for tracking land-use dynamics [58].
 Recent progress in remote sensing technologies and machine learning algorithms has significantly enhanced the ability to monitor agricultural land use.
 High-resolution, multi-spectral data from platforms such as Sentinel-2 provide frequent, detailed observations of croplands at minimal cost.
-By applying advanced machine learning methods, including conventional ensemble learners, e.g, XGBoost (Chen and Guestrin,, 2016) and deep neural networks, e.g, Transformers (Vaswani et al.,, 2017), to these time series images, we can capture crop-specific growth dynamics and distinguish among diverse management practices with high precision.
-These integrated approaches enable automated, scalable mapping of crop types and growth stages, supporting timely, data driven decisions for sustainable agriculture (Turkoglu,, 2023).
-Figure 1: Cumulative Growing Degree Days (GDD) across Switzerland for multiple years.
-Modern deep learning methods (Garnot and Landrieu, (2021), Turkoglu et al., 2021a , Metzger et al., (2021), Tarasiou et al., (2023), Rußwurm and Körner, (2018)) excel in classifying crop types from optical satellite image time series.
-However, these models are typically trained on curated datasets with limited size and/or validated on data collected within a single growing season (Rußwurm et al., (2019), Garnot and Landrieu, (2021), Schneider et al., (2023), Turkoglu et al., 2021a ), neglecting interannual shifts in growth driven by climate variability and thus limiting their applicability across years and regions.
+By applying advanced machine learning methods, including conventional ensemble learners, e.g, XGBoost [8] and deep neural networks, e.g, Transformers [56], to these time series images, we can capture crop-specific growth dynamics and distinguish among diverse management practices with high precision.
+These integrated approaches enable automated, scalable mapping of crop types and growth stages, supporting timely, data driven decisions for sustainable agriculture [50].
+Figure 1: Spatially averaged cumulative growing degree days (cGDD) across Switzerland for multiple years.
+Modern deep learning methods [14, 51, 29, 48, 40] excel in classifying crop types from optical satellite image time series.
+However, these models are typically trained on curated datasets with limited size and/or validated on data collected within a single growing season [39, 42, 46, 14, 51, 35], neglecting interannual shifts in growth driven by climate variability and thus limiting their applicability across years and regions.
 Furthermore, their dependence on current-year labeled data for training precludes real time deployment during the ongoing season.
-As also noted by Roscher et al., (2024), model development should focus on datasets with real-world relevance to further improve accuracy, generalization ability, and real impact on end user applications.
+As also noted by [37], model development should focus on datasets with real-world relevance to further improve accuracy, generalization ability, and real impact on end user applications.
 Lastly, although these approaches may achieve state-of-the-art accuracy, they rarely incorporate uncertainty quantification, leaving predictive confidence unassessed and often uncalibrated.
 Ensuring the models are well-calibrated makes them reliable components of agricultural monitoring systems, where miscalibration could otherwise lead to costly resource misallocations or risk mismanagement.
-Figure 2: NDVI time series of two parcels with the same crop rotation over two years (2021 and 2022), aligned by calendar date (left) and by cumulative growing degree days (GDD) (right).
+Figure 2: NDVI time series of two parcels with the same crop rotation over two years (2021 and 2022), aligned by calendar date (left) and by cumulative growing degree days (cGDD) (right).
 Seen from an ecophysiological point of view, crop development is majorly driven by thermal conditions, since temperature plays a pivotal role in determining growth and phenological development.
@@
 Figure 3: Year-to-year variation of the same agricultural landscape on June 15th from 2021 to 2023.
-In this work, we build upon the state-of-the-art attention-based deep learning architecture proposed by Garnot and Landrieu, (2021), and introduce a simple, model-agnostic sampling strategy that incorporates daily average temperature to reduce redundancy and noise in satellite image time series.
+In this work, we build upon the state-of-the-art attention-based deep learning architecture proposed by Garnot and Landrieu [14], and introduce a simple, model-agnostic sampling strategy that incorporates daily average temperature to reduce redundancy and noise in satellite image time series.
 Deep learning models for crop classification typically rely on subsampled time series, as crop development progresses gradually and many satellite observations are temporally redundant.
 Subsampling is also necessary to reduce sequence length for efficient model training and inference within limited GPU memory.
-To better guide this selection, we replace calendar days with growing degree days (GDD) as the temporal reference and perform uniform subsampling in this thermal time domain.
-This approach emphasizes the crop’s biologically active periods, ensuring that the model focuses on the most informative observations.
+To better guide this sampling, we reparameterize calendar time into thermal time, using cumulative growing degree days (GDD) as our temporal reference, and then perform uniform subsampling in this thermal‐time domain.
+This ensures the model concentrates on the crop’s biologically active periods, focusing on the most informative observations.
 We evaluate our GDD-informed sampling approach on a multi-year, country-wide dataset spanning three Swiss growing seasons by training on one year and testing on the other two.
@@
 2 Related Work
-Deep Learning for Crop Mapping from Satellite Imagery.
-Deep learning has dramatically improved crop‐type classification from satellite time series, outperforming traditional methods based on handcrafted features (Turkoglu,, 2023).
-Early sequence‐modeling approaches used recurrent architectures to capture phenological dynamics.
-Long short‐term memory (LSTM) networks demonstrated strong performance on multi‐temporal inputs (Rußwurm and Körner,, 2017), and the convolutional‐RNN (convRNN) version further boosted both accuracy and efficiency (Turkoglu et al., 2021a, ).
-Attention mechanisms have emerged as a powerful alternative for modeling long‐range temporal dependencies.
-Rußwurm and Körner, (2020) and Garnot et al., (2020) were the first to apply self‐attention to crop classification from Sentinel-2 data and showed improved results.
-Series, (2020) then proposed the lightweight Temporal Attention Encoder (L-TAE), and Garnot and Landrieu, (2021) embedded L-TAE within a U-Net backbone to adaptively weigh temporal features across space and time.
-More recently, fully attention-based Vision Transformer (ViT) adaptations have been explored for crop mapping (Bai et al.,, 2023), and Qin et al., (2024) leveraged the recently proposed Mamba state‐space architecture to improve spatio‐temporal classification performance.
-To incorporate domain knowledge in deep learning architectures, several works introduce structured priors and hierarchical labels.
-Turkoglu et al., 2024b exploited multi‐scale crop class hierarchies within a parameter‐efficient convRNN (i.e, convSTAR (Turkoglu et al., 2021b, )) framework, and prototype‐based supervision has been used to integrate class relationships into metric learning (Sainte Fare Garnot and Landrieu,, 2020).
-Positional encodings tailored to agricultural cycles have also been proposed: (Nyborg et al.,, 2022) introduced thermal-time encoding, using cumulative growing degree days, to normalize phenological shifts across regions.
-This is the most relevant research for this paper.
-Moreover, practical deployment demands handling early-season data and irregular revisit intervals.
-Rußwurm et al., (2023) added an early classification reward loss to LSTM training, improving accuracy in the season when only partial time series are available.
-Metzger et al., (2021) leveraged neural ordinary differential equations (NODE) to process noisy and irregularly sampled time series, and Cai et al., (2023)’s collect–update–distribute design further addresses sampling irregularity.
-Multi-Year Crop Mapping from Satellite Imagery.
-Multi-year crop mapping faces challenges due to temporal domain shifts such as interannual variability in weather, phenological differences, and sensor calibration issues.
-These factors often degrade the performance of models trained on single-season data when applied across multiple years.
-Turkoglu et al., 2024a extended the experiments of Turkoglu et al., 2021a to a multi‐year setting, training and testing on data from two different years, and reporting preliminary results.
-Blickensdörfer et al., (2022) employed a Random Forest classifier using dense Sentinel-2 and Landsat-8 time series, monthly Sentinel-1 composites, and environmental variables to evaluate the relative contributions of optical, radar, and environmental data across three meteorologically distinct years (2017–2019).
-Dubrovin et al., (2024) demonstrated that reconstructing multi-year Sentinel-2 NDVI time series with Fourier-based fitting and applying Gradient Boosting enables improved crop mapping across different years.
-Gao et al., (2023) investigated ways to optimize the use of multi-year samples in a within-season crop classification model.
-Qi et al., (2024) presented a Google Earth Engine workflow that fuses field-sample data with multi-year Sentinel-2 time series to map cropping systems on smallholder farms, achieving consistently robust results across diverse landscapes.
-Li et al., (2023) benchmarked deep learning segmentation models (e.g., U-Net and SegFormer (Xie et al.,, 2021)) for crop mapping across multiple years, while Sykas et al., (2022) evaluated convLSTM and convSTAR (Turkoglu et al., 2021b, ) networks in a similar multi-year context.
-Very recently, Khallaghi et al., (2025) shows improved results for cross-year crop mapping combining a U-Net with photometric augmentation, Tversky-Focal loss, and Monte Carlo (MC) Dropout.
-Uncertainty Modeling in Deep Learning.
-Predictive uncertainty is commonly split into aleatoric uncertainty, which stems from inherent data variability such as class overlap or sensor noise, and epistemic uncertainty, which reflects model weight uncertainty due to insufficient coverage of the input space (Kendall and Gal,, 2017).
-Exact Bayesian inference in deep networks is generally intractable, leading to approximate methods like variational inference (e.g., Bayes by Backprop,  Blundell et al., 2015), Markov Chain Monte Carlo techniques such as Stochastic Gradient Langevin Dynamics (Welling and Teh,, 2011).
-However, these approaches often struggle to scale to large architectures (Gustafsson et al.,, 2019).
-Probabilistic ensembles, particularly deep ensembles, currently represent the gold standard for epistemic uncertainty estimation in deep networks, providing well‐calibrated and robust predictions by aggregating outputs from multiple independently trained models (Lakshminarayanan et al.,, 2017).
-However, the need to train and store several large networks incurs substantial computational and memory overhead.
-To address this, implicit ensemble techniques, such as Monte Carlo (MC) Dropout (Gal and Ghahramani,, 2016), Snapshot Ensembles (Huang et al.,, 2017), Masksembles (Durasov et al.,, 2021), and FiLM-Ensemble (Turkoglu et al.,, 2022), approximate ensemble diversity via stochastic weight perturbations or dynamic weight combinations, significantly reducing resource requirements.
-Despite these gains in efficiency, implicit methods generally underperform full ensembles in both predictive accuracy and uncertainty calibration.
-More recently, LoRA-Ensemble has been introduced as a parameter‐efficient alternative for transformer‐based architectures, achieving uncertainty estimates competitive with deep ensembles at a fraction of the cost (Mühlematter et al.,, 2024).
-Uncertainty Modeling in Satellite Imagery and Crop Mapping
-Uncertainty quantification has become a standard component of satellite-based land-cover and crop-mapping workflows, as it provides essential information for risk-aware decision-making in agricultural management and ecosystem monitoring.
-Generating per-pixel or per-field uncertainty maps alongside deterministic predictions enables practitioners to identify low-confidence areas and tailor downstream analyses accordingly.
-For example, Lang et al., (2022) employs a deep ensemble of convolutional neural networks to estimate global canopy height from sparse LiDAR samples and multi-sensor satellite imagery, producing uncertainty intervals that reflect both model and data variability.
-Becker et al., (2023) extends this ensemble approach to country-scale forest structure estimation at 10m resolution, demonstrating that dense uncertainty fields help flag regions where auxiliary field data are most needed.
-In the context of crop yield and type mapping, several recent studies have adopted stochastic inference techniques.
-Tian et al., (2024) introduces an attention-guided multi-level crop network for winter wheat yield estimation at the county scale, applying MC Dropout during inference to derive spatially explicit uncertainty maps of yield forecasts.
-Similarly, Khallaghi et al., (2025) leverages MC Dropout in a U-Net segmentation framework to quantify per-pixel uncertainty in annual cropland masks across multiple years in Ghana, using the resulting confidence estimates to discard unreliable predictions.
-Liu et al., (2024) integrates MC Dropout in a ViT–ResNet feature encoder to produce field-level uncertainty scores, filtering out low-confidence labels to improve overall mapping precision.
-Beyond dropout-based methods, Baumert et al., (2024) proposes a probabilistic crop-type mapping method in which an ensemble of classifiers, each trained with perturbed inputs and hyperparameters on Sentinel-1, Sentinel-2, and Landsat-8 data, generates multiple crop maps over Germany; per-pixel frequency counts across the ensemble yield robust probability distributions and uncertainty metrics.
-Maleki et al., (2024) refines the 30m cropland data by harnessing its built-in Random Forest confidence layer: low-certainty regions are iteratively reclassified using targeted ancillary datasets, resulting in a more accurate and self-aware mapping product.
-Collectively, these approaches demonstrate the critical role of uncertainty modeling in improving the reliability and interpretability of satellite-driven crop and canopy mapping systems.
+Deep learning has revolutionized crop‐type classification from satellite time series.
+Early sequence models based on LSTM [38] and convolutional‐RNN architectures [51] harnessed phenological dynamics, while attention‐based approaches such as self‐attention [41, 15], the lightweight Temporal Attention Encoder (L‐TAE) [44], and U‐Net–integrated attention [14] improved long‐range dependencies.
+Vision Transformer adaptations [2] and state‐space models like Mamba [33] further enhanced spatio‐temporal representations.
+Integrating domain knowledge, convSTAR [55] and prototype‐based supervision [45] has been used to incorporate class relationships.
+The most relevant method for crop type mapping is thermal-time positional encoding [31], which aligns phenological stages across regions using cumulative growing degree days.
+To handle irregular sampling and early‐season needs, methods incorporating reward‐based early classification [43], neural ODEs [29], and adaptive sampling frameworks [7] have been developed.
+Generalizing across seasons is hampered by interannual phenological shifts and sensor inconsistencies.
+Multi‐year evaluations [54, 5] and methods such as Fourier‐based time series reconstruction with gradient boosting [9], Google Earth Engine workflows [32], and photometric augmentation in U‐Net models [22] have shown promising improvements.
+Uncertainty estimation enhances trust in crop maps.
+Deep ensembles [23] remain the standard, with efficient approximations including MC Dropout [12], Snapshot Ensembles [20], Masksembles [10], FiLM‐Ensemble [53], and LoRA‐Ensemble [19].
+In satellite‐based crop monitoring, stochastic inference quantifies per‐pixel and per‐field uncertainty, guiding data collection and decision‐making.
+MC Dropout has been applied to winter wheat yield forecasting [49], annual cropland mask segmentation [22], and ViT–ResNet feature encoding for field‐level confidence scoring [27].
+Probabilistic deep ensembles trained on multi‐sensor inputs produce robust uncertainty distributions, as demonstrated in German crop‐type mapping [3].
+Integrating these uncertainty estimates with deterministic predictions flags low‐confidence areas, enabling targeted ground truthing and more reliable operational deployment of crop‐type mapping systems.
+Refer to Appendix C for an extended version of the Related Work section.
 3 Method
-Our method builds on the U-Net with Temporal Attention Encoder (U-TAE) (Ronneberger et al.,, 2015; Garnot and Landrieu,, 2021), a state-of-the-art deep learning model for dense segmentation of satellite image time series.
+Formally, the objective is to predict a crop type map $\bm{y}\in{\rm I\!R}^{H\times W\times C}$ from a sequence of input images $\bm{X}=\{\bm{x}_{1},\bm{x}_{2},..,\bm{x}_{L}\}\in{\rm I\!R}^{H\times W\times B}$.
+$H$ and $W$ are the height and width of the input images, respectively, $B$ is the number of input bands, $L$ is the number of time stamps in the input sequence that can differ from sample to sample, and $C$ is the number of crop types.
+Our method builds on the U-Net with Temporal Attention Encoder (U-TAE) [36, 14], a state-of-the-art deep learning model for dense segmentation of satellite image time series.
 Prior to presenting our sampling approach, we examine U-TAE on a standardized benchmark by adjusting sequence lengths and positional encodings to identify the minimal temporal resolution and ordering information needed for high classification accuracy.
@@
 Figure 4: Learning curves of the U-TAE model with different position encodings (PE) on the PASTIS dataset, showing training loss, training accuracy, validation loss, and validation accuracy respectively.
-3.1 U-Net with Temporal Attention Encoder (U-TAE)
-The U-TAE model is a spatio-temporal neural network designed to extract dense and temporally-aware features from satellite image time series.
-It consists of a shared 2D convolutional encoder applied to each timestamp independently, followed by a temporal self-attention module that aggregates the encoded features across time.
-This attention module uses positional encoding to preserve temporal ordering and applies scaled dot-product attention to dynamically weigh each timestamp based on its relevance.
-The aggregated multi-temporal representation is then decoded through a U-Net-style decoder with skip connections, enabling high-resolution semantic segmentation.
-The architecture can handle variable-length input sequences and produce temporally adaptive representations of land cover.
-Refer to Garnot and Landrieu, (2021) for more details.
+3.1 U-Net w/ Temporal Attention Encoder (U-TAE)
+The U-TAE model fuses spatial feature extraction with temporal attention to segment land cover in satellite image time series $\{\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{L}\}$.
+Each frame $\mathbf{x}_{i}$ is passed through a shared 2D convolutional encoder, producing per-frame feature maps.
+These maps are then aggregated by a self-attention block that employs positional encodings and scaled dot-product attention to weight timestamps according to their phenological relevance.
+The resulting multi-temporal embedding is fed into a U-Net–style decoder with skip connections, yielding high-resolution, temporally adaptive segmentation masks.
+U-TAE naturally handles sequences of varying length and is optimized via pixel-wise cross-entropy loss against the ground-truth labels.
+For training details, refer to the Appendix A.
 3.2 Preliminary Analysis with U-TAE
@@
 There is currently no comprehensive study establishing the exact number of time steps required to classify crops reliably using U-TAE or other architectures.
-Existing works vary widely, for example, Garnot and Landrieu, (2021) uses up to 61 observations, while Turkoglu et al., 2021a employs 71 timestamps.
+Existing works vary widely, for example, Garnot and Landrieu [14] uses up to 61 observations, while Turkoglu et al.
+[51] employs 71 timestamps.
 In principle, Sentinel-2 can provide an observation every 2 to 3 days over Switzerland, making very long sequences feasible.
-However, such long sequences are often redundant and noisy: Metzger et al., (2021) reports that roughly 50% of observations can be unusable due to cloud cover.
+However, such long sequences are often redundant and noisy: Metzger et al.
+[29] reports that roughly 50% of observations can be unusable due to cloud cover.
 Moreover, crops may develop relatively slowly, e.g.
@@
 To maintain a reasonable batch size during training, it is essential to limit the number of timestamps.
-Note that larger batches reduce the variance in the estimation of the gradient, resulting in a stronger and more stable training signal (Räisä et al.,, 2024).
-This is particularly advantageous for highly imbalanced datasets: choosing a batch size several times the minority-to-majority class ratio ensures that each minibatch contains sufficient minority-class examples to learn meaningful features (Google Developers,, 2025).
-To ensure our evaluation remains unbiased, we base our initial analysis on the well-established PASTIS benchmark, a concise and reproducible toy crop classification dataset introduced alongside U-TAE by Garnot et al.
-(2021).
+Note that larger batches reduce the variance in the estimation of the gradient, resulting in a stronger and more stable training signal [34].
+This is particularly advantageous for highly imbalanced datasets: choosing a batch size several times the minority-to-majority class ratio ensures that each minibatch contains sufficient minority-class examples to learn meaningful features [17].
+To ensure our evaluation remains unbiased, we base our initial analysis on the commonly used, well-established PASTIS benchmark, a concise and reproducible toy crop classification dataset introduced alongside U-TAE by [14].
 By mirroring SwissCrop’s multi-temporal structure while remaining fully independent, PASTIS enables rapid iteration and guarantees that insights from our analysis directly inform algorithm design without the dataset itself biasing the method.
 In this dataset each time series contains on average 48 Sentinel-2 observations (up to 61).
-We uniformly subsampled each sequence to 24 time steps, and trained with a batch size of 4 (see Garnot and Landrieu, (2021) for full training details).
+We uniformly subsampled each sequence to 24 time steps, and trained with a batch size of 4 (see [14] for full training details).
 As shown in Table 1, this simple uniform subsampling yields only a 0.2% absolute drop in overall accuracy, despite no selective filtering of cloudy or low-information dates.
 Crucially, GPU memory usage is cut in half, and the number of multiply–accumulate operations (MACs) per forward pass with a single batch is reduced around 50%, dramatically lowering computational cost and freeing capacity for larger batches or faster training and inference.
-Accordingly, we set the temporal length to 24 for the experiments.
-Table 1: Performance comparison for varying temporal length on the Pastis dataset.
-Results are averaged over a 5-fold experimental setting of the Pastis dataset.
-Refer to Garnot and Landrieu, (2021) for details.
-M stands for GPU memory.
+Accordingly, we set the temporal length $T=24$ for the experiments.
+Table 1: Performance comparison for varying temporal length on the Pastis dataset. Results are averaged over a 5-fold experimental setting. Refer to [14] for details. M stands for GPU memory.
 $T_{min}$ $T_{max}$ $T_{mean}$ M (Gb) MACs (B) Acc (%)
@@
   • linearPE: Positions are encoded as sequential integers (1, 2, 3, …).
-  • calendarPE: Positions are encoded as the actual calendar dates of each observation (1, …, 365).
+  • calendarPE: Positions are encoded as the actual calendar dates of each observation, e.g., (2, 5,…, 365).
 As shown in Figure 4, the noPE model underperforms in both training and validation metrics, confirming that some ordering information is necessary.
@@
 This result led us to conclude that using thermal time directly as a positional encoding is unlikely to be very effective, as we show in our experiments.
-In summary, our preliminary findings reveal that (i) substantial temporal subsampling incurs only a marginal accuracy loss while reducing both GPU memory usage and computational cost drastically, and (ii) U-TAE relies chiefly on the relative ordering of observations, whereas absolute timestamps offer limited benefit.
+In summary, our preliminary findings reveal that (i) substantial temporal subsampling incurs only a marginal accuracy loss while drastically reducing both GPU memory usage and computational cost, and (ii) U-TAE relies chiefly on the relative ordering of observations, whereas absolute timestamps offer limited benefit.
 These insights suggest that, rather than encoding raw dates, we should focus on selecting the most informative, developmentally relevant observations.
@@
 3.3 T3S: Thermal-Time-based Temporal Sampling
-Satellite-based crop monitoring approaches often rely on fixed-interval calendar-day sampling to construct input sequences.
-However, this overlooks the ecophysiological processes of plant growth such as their dependence on temperature.
-As crop development is tightly coupled with temperature, we propose a simple yet effective model-agnostic sampling method, Thermal-Time-based Temporal Sampling (T3S), which replaces fixed calendar intervals with intervals defined in thermal time, using Growing Degree Days (GDD) as the temporal reference.
-GDD provides a cumulative measure of heat accumulation, reflecting the physiological development of crops more faithfully than absolute time. For a given day $d$, GDD is computed as:
-$\mathrm{GDD}_{d}=\max\left(0,\frac{T_{\text{max},d}+T_{\text{min},d}}{2}-T_{\text{base}}\right),$ (1)
-where $T_{\text{max},d}$ and $T_{\text{min},d}$ are the daily maximum and minimum temperatures, respectively, and $T_{\text{base}}$ is a base threshold below which crop development is considered negligible.
+Satellite-based crop monitoring often relies on fixed-interval calendar‑day sampling to construct input sequences, overlooking temperature‑driven plant ecophysiology.
+Since crop development is tightly coupled with accumulated heat, we introduce a simple yet effective, model‑agnostic sampling method, T3S, which reparametrizes calendar time into thermal time, replacing fixed calendar intervals with intervals defined by growing degree days (GDD) as the temporal reference, and then performs uniform subsampling in this thermal‑time domain.
+GDD provides a cumulative measure of heat accumulation, reflecting the physiological development of crops more faithfully than absolute time. For a given day $i$, GDD is computed as
+$\mathrm{GDD}_{i}=\max\left(0,\frac{T_{\text{max},i}+T_{\text{min},i}}{2}-T_{\text{base}}\right),$ (1)
+where $T_{\text{max},i}$ and $T_{\text{min},i}$ are the daily maximum and minimum temperatures for day $i$, respectively, and $T_{\text{base}}$ is a base threshold below which crop development is considered negligible.
 In our experiments, we use $T_{\text{base}}=0^{\circ}$, a commonly used value for temperate climate crops.
-We then compute the cumulative GDD over time:
-$\mathrm{GDD}_{D}=\sum_{d=1}^{D}\mathrm{GDD}_{d}.$ (2)
-To construct a time series of $L$ observations, we partition the full GDD range $\left[\mathrm{GDD}{\min},\mathrm{GDD}{\max}\right]$ into $L$ uniform intervals.
+We then compute the cumulative GDD (cGDD) over time for day d:
+$\mathrm{cGDD}_{d}=\sum_{i=1}^{d}\mathrm{GDD}_{i}.$ (2)
+To construct a time series of $T(<L)$ observations, we partition the full cGDD range $\left[\mathrm{cGDD}{\min},\mathrm{cGDD}{\max}\right]$ into $T$ uniform intervals.
 Within each interval, we select the least cloudy satellite observation, prioritizing data quality.
@@
 The T3S algorithm is outlined in Algorithm 1.
-It takes as input the full set of timestamps, the cumulative thermal time values, i.e, GDD, the satellite data cube, a cloud mask for each observation, and the desired number of samples $L$.
+It takes as input the full set of timestamps $\{1,2,...,L\}$, where $L$ is the initial time length, the satellite data $\bm{X}\in{\rm I\!R}^{L\times H\times W\times B}$, the cumulative thermal time values, i.e, $cGDD$, a cloud mask for each observation, and the desired number of samples $T$.
 For each thermal interval, it identifies candidate observations and selects the one with the fewest cloud-contaminated pixels.
@@
 Lastly, T3S is agnostic to model architecture and can be seamlessly integrated into any time series-based crop monitoring pipeline.
-Figure 5: The first row shows conventional uniform sampling; the second row shows uniform sampling where the least cloudy observation is chosen in each interval (referred to as “Deformable” baseline); the third row shows the proposed Thermal‐Time‐based Temporal Sampling (T3S) approach.
-In T3S, each square box represents the temperature sum accumulated since the previous one, illustrating intervals with equal thermal time.
+Figure 5: Method overview.
+The first row shows conventional uniform sampling; the second row shows uniform sampling where the least cloudy observation is chosen in each interval (referred to as “Deformable” baseline); the third row shows the proposed Thermal Time‐based Temporal Sampling (T3S).
+In T3S, each square box represents the temperature sum accumulated since the previous one, illustrating intervals with equal cumulative GDD.
 Algorithm 1 T3S
-1:function T3S(timeStamps, GDD, data, cloudMask, $L$)
-2:     $t_{\min}\leftarrow\min(\text{GDD})$
-3:     $t_{\max}\leftarrow\max(\text{GDD})$
-4:     $\Delta\leftarrow(\,t_{\max}-t_{\min}\,)/L$   #$L$ is target temporal length
+1:function T3S(timeStamps, $\bm{X}$, cGDD, cloudMask, $T$)
+2:     $t_{\min}\leftarrow\min(\text{cGDD})$     # $\bm{X}$’s shape is [L,H,W,B]
+3:     $t_{\max}\leftarrow\max(\text{cGDD})$
+4:     $\Delta\leftarrow(\,t_{\max}-t_{\min}\,)/T$   #$T$ is target temp. length
 5:     $selected\leftarrow[\ ]$
 6:     $current\leftarrow t_{\min}$
-7:     while $|selected|<L$ and $current<t_{\max}$ do
+7:     while $|selected|<T$ and $current<t_{\max}$ do
 8:         $next\leftarrow current+\Delta$
-9:         $window\leftarrow\{\,i\mid current\leq\text{GDD}[i]<next\}$
+9:         $window\leftarrow\{\,i\mid current\leq\text{cGDD}[i]<next\}$
 10:         if $window\neq\emptyset$ then
-11:              $best\leftarrow\displaystyle\arg\max_{i\in window}\sum(\text{cloudMask}[i]=0)$
+11:              $best\leftarrow\displaystyle\operatorname*{arg\,max}_{i\in window}\sum(\text{cloudMask}[i]=0)$
 12:              append $best$ to $selected$
@@
 16:     sort($selected$)
-17:     return data[:, selected, …]
+17:     return $\bm{X}$[selected, …]  # the shape is [T,H,W,B]
 18:end function
-3.4 Training Details
-The model was trained for 100 epochs using a batch size of 8 and a fixed temporal length of 24.
-We employed the Adam optimizer with an initial learning rate of 0.001, scheduled by a OneCycleLR policy with a maximum learning rate of 0.01 and a warm-up phase covering 30% of the total training steps.
-The U-TAE architecture is configured with an encoder having channel widths $[128,128,128,256]$, a symmetric decoder with widths $[32,64,128,256]$ and strided convolutions with kernel size $k=4$, stride $s=2$, and padding $p=1$.
-The temporal attention module employs $n_{\text{head}}=16$ heads and a model dimension of $d_{\text{model}}=256$.
-Input features are standardized to zero mean and unit variance on a per-channel basis.
-We implemented the model in PyTorch and trained it on a single NVIDIA RTX 4090 GPU over roughly three days.
-The full codebase and pretrained checkpoints will be publicly available on GitHub.
 4 Dataset
-In this work, we work with SwissCrop 2021, 2022, 2023, a three-year crop dataset covering the entirety of Switzerland from 2021 to 2023.
-SwissCrop combines Sentinel-2 Level-2A bottom-of-atmosphere multi-spectral imagery with annual crop‐type labels for each field.
+We work with SwissCrop 2021, 2022, 2023, a three-year crop dataset covering the entirety of Switzerland (see Figure 10 in the appendix).
+SwissCrop combines Sentinel-2 Level-2A bottom-of-atmosphere multi-spectral imagery, averaging 110 timestamps per season, with annual crop‐type labels for each field.
 The dataset includes 50 distinct crop types, each representing the primary crop grown in a field during a given season.
-These labels where provided by farmers to the cantons of Switzerland according to a data model provided by the Federal Office for Agriculture  [^1] https://www.blw.admin.ch/de/landwirtschaftliche-kulturflaechen and joined together by the Swiss Centre for Agricultural Research, Agroscope.
-Unlike many existing datasets, e.g.
-Pastis dataset (Garnot and Landrieu,, 2021), SwissCrop faithfully reflects the pronounced class imbalance found in real‐world farming: a long‐tail distribution in which a handful of major crops dominate while the majority occur only sporadically (cf.
-Figure 6).
-[…]
-To complement these remote sensing observations, we incorporate high-resolution gridded climate data from MeteoSwiss, which supplies near-surface temperature fields (minimum, maximum, and average) on 1 km, 2 km, and 5 km grids (Federal Office of Meteorology and Climatology, MeteoSwiss,, 2024).
-In our study, we use the 1 km daily minimum (TminD) and maximum (TmaxD) temperature series to compute growing degree days (GDD; refer to Section 3.3).
-We make the full SwissCrop dataset publicly available to the research community.
-We hope this resource will facilitate further advances in large-scale, cross-year crop classification, phenology modeling, and earth observation research.
-Figure 6: SwissCrop dataset label distribution for the year 2021. Note the logarithmic scale of the y-axis.
-Figure 7: Region of interest (top left), the SwissCrop dataset (bottom left) and its coverage by data cubes (center), with an example of a single analysis-ready cube (right).
-5 Experiment & Results
-We evaluate the proposed T3S against four very relevant state-of-the-art baselines using a leave‐one‐year‐out scheme over the 2021–2023 SwissCrop dataset.
+Unlike many existing datasets, e.g., commonly used Pastis benchmark [14], SwissCrop faithfully reflects the pronounced class imbalance found in real‐world farming: a long‐tail distribution in which a handful of major crops dominate while the majority occur only sporadically (see Figure 11 in the appendix for the class distribution of SwissCrop).
+To drive phenology-aware sampling, we complement the imagery with daily temperature data (originally at 1 km resolution and resampled to 10 m to match Sentinel-2 imagery) from MeteoSwiss to compute growing degree days (GDD).
+The full SwissCrop dataset, including imagery, labels, and temperature series, is publicly released to facilitate reproducible research in large-scale crop mapping, phenology modeling, and earth observation research.
+For more details, refer to the Appendix B.
+5 Experiment & Discussion
+We evaluate the proposed T3S against four relevant state-of-the-art baselines using a leave‐one‐year‐out scheme over the 2021–2023 SwissCrop dataset.
 In each of six cross‐year folds, two seasons are held out for testing while the remaining one season is used for training.
-To further assess robustness, we conduct a low‐data experiment using only 10% of the labels, and an early‐season study by truncating each test timeseries at the 50th and 75th time percentile (end of June and end of September respectively).
+To further assess robustness, we conduct a low‐data experiment using only 10% of the labels, and an early‐season study by truncating each test time series at the 50th and 75th time percentile (end of June and end of September, respectively).
 We assess model predictive performance using overall accuracy, the fraction of pixels correctly labeled; intersection‐over‐union (IoU), the ratio of the area of overlap between predicted and ground‐truth pixels to the area of their union; and mean intersection‐over‐union (mIoU), averaged across classes.
-Calibration is measured by the Expected Calibration Error (ECE), which measures how far predicted confidences deviate from observed error rates, i.e., perfect calibration occurs when the estimated uncertainties match the actual likelihood of a misclassification.
-We also report Negative Log‐Likelihood (NLL), the mean negative log‐probability of the true class, and the Brier Score, the mean squared error between predicted probabilities and one‐shot labels.
+Calibration is measured by the expected calibration error (ECE), which measures how far predicted confidences deviate from observed error rates, i.e., perfect calibration occurs when estimated uncertainties match the actual likelihood of misclassification.
+We also report negative log‐likelihood (NLL), the mean negative log‐probability of the true class, and Brier score, the mean squared error between predicted probabilities and one‐shot labels.
 Table 2: Performance comparison across six evaluation settings.
@@
 Train Test Method Accuracy % ($\uparrow$) mIoU % ($\uparrow$) IoU % ($\uparrow$) ECE ($\downarrow$) NLL ($\downarrow$) Brier Score ($\downarrow$)
-2021 2022 U-TAE (Garnot and Landrieu,, 2021) 70.1 17.2 53.9 0.053 0.943 0.412
-+MC-Dropout (Gal and Ghahramani,, 2016) 71.3 17.4 55.4 0.032 0.911 0.402
-+Thermal-PE (Nyborg et al.,, 2022) 72.5 18.6 56.9 0.045 0.854 0.385
+2021 2022 U-TAE [14] 70.1 17.2 53.9 0.053 0.943 0.412
++MC-Dropout [12] 71.3 17.4 55.4 0.032 0.911 0.402
++Thermal-PE [31] 72.5 18.6 56.9 0.045 0.854 0.385
 +Deformable Sampling 73.8 20.5 58.5 0.044 0.836 0.371
 T3S (ours) 76.0 21.6 61.3 0.004 0.720 0.329
-2021 2023 U-TAE (Garnot and Landrieu,, 2021) 71.4 16.7 55.6 0.052 0.947 0.399
-+MC-Dropout (Gal and Ghahramani,, 2016) 72.5 17.0 56.8 0.037 0.918 0.394
-+Thermal-PE (Nyborg et al.,, 2022) 71.3 15.5 55.4 0.035 0.925 0.401
+2021 2023 U-TAE [14] 71.4 16.7 55.6 0.052 0.947 0.399
++MC-Dropout [12] 72.5 17.0 56.8 0.037 0.918 0.394
++Thermal-PE [31] 71.3 15.5 55.4 0.035 0.925 0.401
 +Deformable Sampling 74.6 18.9 59.6 0.032 0.822 0.359
 T3S (ours) 77.6 21.0 63.5 0.005 0.691 0.307
-2022 2021 U-TAE (Garnot and Landrieu,, 2021) 70.8 16.2 54.8 0.035 0.936 0.408
-+MC-Dropout (Gal and Ghahramani,, 2016) 71.2 16.2 55.3 0.024 0.913 0.399
-+Thermal-PE (Nyborg et al.,, 2022) 72.8 18.2 57.2 0.077 1.022 0.403
+2022 2021 U-TAE [14] 70.8 16.2 54.8 0.035 0.936 0.408
++MC-Dropout [12] 71.2 16.2 55.3 0.024 0.913 0.399
++Thermal-PE [31] 72.8 18.2 57.2 0.077 1.022 0.403
 +Deformable Sampling 76.8 20.4 62.3 0.043 0.833 0.339
 T3S (ours) 77.3 20.1 63.0 0.013 0.697 0.310
-2022 2023 U-TAE (Garnot and Landrieu,, 2021) 72.1 17.4 56.4 0.041 0.875 0.389
-+MC-Dropout (Gal and Ghahramani,, 2016) 72.8 17.2 57.2 0.028 0.853 0.381
-+Thermal-PE (Nyborg et al.,, 2022) 63.9 14.2 46.9 0.123 1.374 0.525
+2022 2023 U-TAE [14] 72.1 17.4 56.4 0.041 0.875 0.389
++MC-Dropout [12] 72.8 17.2 57.2 0.028 0.853 0.381
++Thermal-PE [31] 63.9 14.2 46.9 0.123 1.374 0.525
 +Deformable Sampling 76.0 20.8 61.4 0.073 0.936 0.363
 T3S (ours) 76.8 20.9 62.4 0.019 0.708 0.318
-2023 2021 U-TAE (Garnot and Landrieu,, 2021) 71.6 17.6 55.8 0.066 0.913 0.398
-+MC-Dropout (Gal and Ghahramani,, 2016) 71.6 17.6 55.7 0.040 0.889 0.391
-+Thermal-PE (Nyborg et al.,, 2022) 71.8 16.3 56.0 0.098 0.987 0.413
+2023 2021 U-TAE [14] 71.6 17.6 55.8 0.066 0.913 0.398
++MC-Dropout [12] 71.6 17.6 55.7 0.040 0.889 0.391
++Thermal-PE [31] 71.8 16.3 56.0 0.098 0.987 0.413
 +Deformable Sampling 75.3 20.5 60.4 0.105 0.951 0.379
 T3S (ours) 77.1 23.2 62.8 0.019 0.686 0.307
-2023 2022 U-TAE (Garnot and Landrieu,, 2021) 71.2 18.4 55.3 0.059 0.957 0.410
-+MC-Dropout (Gal and Ghahramani,, 2016) 72.0 18.6 56.2 0.043 0.929 0.400
-+Thermal-PE (Nyborg et al.,, 2022) 75.3 19.6 60.4 0.085 0.929 0.366
+2023 2022 U-TAE [14] 71.2 18.4 55.3 0.059 0.957 0.410
++MC-Dropout [12] 72.0 18.6 56.2 0.043 0.929 0.400
++Thermal-PE [31] 75.3 19.6 60.4 0.085 0.929 0.366
 +Deformable Sampling 75.8 21.9 61.0 0.081 0.827 0.355
 T3S (ours) 76.9 22.3 62.5 0.006 0.687 0.311
-Average U-TAE (Garnot and Landrieu,, 2021) $71.2\pm 0.6$ $17.3\pm 0.7$ $55.3\pm 0.8$ $0.051\pm 0.012$ $0.929\pm 0.028$ $0.403\pm 0.008$
-+MC-Dropout (Gal and Ghahramani,, 2016) $71.9\pm 0.6$ $17.3\pm 0.7$ $56.1\pm 0.7$ $\underline{0.034}\pm 0.007$ $0.902\pm 0.029$ $0.395\pm 0.007$
-+Thermal-PE (Nyborg et al.,, 2022) $71.3\pm 3.5$ $17.1\pm 1.9$ $55.5\pm 4.1$ $0.077\pm 0.029$ $1.015\pm 0.156$ $0.416\pm 0.053$
+Average U-TAE [14] $71.2\pm 0.6$ $17.3\pm 0.7$ $55.3\pm 0.8$ $0.051\pm 0.012$ $0.929\pm 0.028$ $0.403\pm 0.008$
++MC-Dropout [12] $71.9\pm 0.6$ $17.3\pm 0.7$ $56.1\pm 0.7$ $\underline{0.034}\pm 0.007$ $0.902\pm 0.029$ $0.395\pm 0.007$
++Thermal-PE [31] $71.3\pm 3.5$ $17.1\pm 1.9$ $55.5\pm 4.1$ $0.077\pm 0.029$ $1.015\pm 0.156$ $0.416\pm 0.053$
 +Deformable Sampling $\underline{75.4}\pm 1.0$ $\underline{20.5}\pm 0.9$ $\underline{60.5}\pm 1.2$ $0.063\pm 0.025$ $\underline{0.868}\pm 0.045$ $\underline{0.361}\pm 0.014$
 T3S (ours) $\textbf{77.0}\pm 0.5$ $\textbf{21.5}\pm 1.0$ $\textbf{62.6}\pm 0.7$ $\textbf{0.011}\pm 0.006$ $\textbf{0.698}\pm 0.012$ $\textbf{0.314}\pm 0.008$
-Figure 8: Class-wise accuracy difference between the proposed T3S and U-TAE.
+Figure 6: Class-wise accuracy difference between T3S and U-TAE.
 Averaged over six experimental folds.
@@
 U-TAE is a U‐Net combined with a temporal attention encoder (i.e., self-attention layer) that ingests images at fixed calendar dates and uses sinusoidal day‐of‐year positional encodings; it serves as our primary reference.
-We combine U-TAE with MC Dropout (Gal and Ghahramani,, 2016): U-TAE+MC‐Dropout retains the same sampling grid but activates dropout layers at test time to perform Monte Carlo sampling, thereby approximating a Bayesian posterior and yielding pixel‐level uncertainty estimates without changing the input dates.
-Note that Monte Carlo Dropout remains one of the most popular and theoretically grounded methods for uncertainty calibration in deep learning (Kendall and Gal,, 2017).
+We combine U-TAE with MC Dropout [12]: U-TAE+MC‐Dropout retains the same sampling grid but activates dropout layers at test time to perform Monte Carlo sampling, thereby approximating a Bayesian posterior and yielding pixel‐level uncertainty estimates without changing the input dates.
+Note that Monte Carlo Dropout remains one of the most popular and theoretically grounded methods for uncertainty calibration in deep learning [21].
 Empirically, the number of ensemble members is set to 5, and the drop rate is set to 0.2.
-U-TAE+Thermal-PE follows Nyborg et al., (2022) and replaces the sinusoidal calendrical encodings with absolute thermal‐time (Growing Degree Days) embeddings, giving the network explicit phenological context while still sampling on calendar dates.
+U-TAE+Thermal-PE follows Nyborg et al.
+[31] and replaces the sinusoidal calendrical encodings with absolute thermal‐time (Growing Degree Days) embeddings, giving the network explicit phenological context while still sampling on calendar dates.
 U-TAE+Deformable Sampling follows the same fixed‐interval calendar schedule (e.g.
@@
 Crucially, T3S also yields much better‐calibrated outputs.
-T3S reduces ECE by roughly fourfold compared to U-TAE, approaching perfect calibration, and substantially lowers both Negative Log‐Likelihood and Brier Score, indicating sharper and more reliable uncertainty estimates.
+T3S reduces ECE by roughly fourfold compared to U-TAE, approaching perfect calibration, and substantially lowers both NLL and Brier score, indicating sharper and more reliable uncertainty estimates.
 By comparison, U-TAE+MC‐Dropout achieves the second‐best calibration, modestly boosting accuracy and reducing ECE relative to vanilla U-TAE but still falling short of T3S.
@@
 In the first example, both models misclassify the highlighted region, but U-TAE remains overconfident in its incorrect label while T3S appropriately elevates uncertainty around the error.
-In the second example, T3S confines its errors to the fuzzy edges of the highlighted field, which is understandable given that farmers hand-draw plot boundaries and 10m pixels often cover two crop types, and its uncertainty map shows a pronounced hotspot exactly along those misclassified borders.
+In the second example, T3S confines its errors to the fuzzy edges of the highlighted field.
+This is understandable given that farmers hand-draw plot boundaries and 10m pixels often cover two crop types, and the T3S uncertainty map shows a pronounced hotspot exactly along those misclassified borders.
 In contrast, U-TAE not only mislabels the entire field but also produces a diffuse, misaligned uncertainty map that neither highlights its own errors nor respects the actual field geometry.
@@
 These examples illustrate that T3S not only reduces classification errors but also delivers well-calibrated uncertainty estimates that expose both model and dataset inconsistencies.
-Low Data Regime.
-In many agricultural settings, annotated crop labels are scarce or expensive to obtain, which makes low data performance a critical requirement for operational viability.
-To assess robustness under label scarcity, we conducted a six-fold experiment using only 10 % of the training data, randomly sampled once and held constant across methods, and evaluated both T3S and U-TAE on the full test set.
-As shown in Figure 9, both models incur drops in accuracy and increases in Expected Calibration Error when labels are limited, but U-TAE’s performance deteriorates far more rapidly.
-Remarkably, T3S trained on just 10 % of the data still outperforms U-TAE trained on the full dataset, and its calibration remains substantially more reliable.
-This advantage stems from the incorporation of a meaningful inductive bias, namely the thermal time prior, which encodes domain knowledge about crop development into the model architecture.
-By constraining the hypothesis space to reflect biologically plausible growth patterns, T3S requires fewer examples to learn robust representations, making it far more data efficient and noise resilient, for example, against cloud cover and sensor artifacts, than models trained purely in a data-driven fashion.
-These results demonstrate that thermal time inductive bias confers strong resilience to small data scenarios, underscoring T3S’s practicality for real-world applications where labels are limited.
-See Section 6 for more discussion.
 Early‐Season Classification.
@@
 To evaluate T3S in this context, we truncate the test set time series at two percentiles, 75 percent (end of September) and 50 percent (end of June), and perform a six fold experiment in which both T3S and U-TAE are trained on full-season data and tested on the truncated series.
-Figure 10 shows that both models experience a decline in in-season performance, measured by predictive accuracy and uncertainty estimation, when using partial time series.
+Figure 7 shows that both models experience a decline in in-season performance when using partial time series.
 At the 75 percent cutoff, U-TAE’s performance drops sharply, whereas T3S remains virtually unchanged.
@@
 These results indicate that T3S not only improves in-season classification accuracy and uncertainty quantification but also underscores its practical suitability for operational in-season crop monitoring and decision-support applications, where timely and reliable predictions are critical.
-Figure 9: Low data regime performance showing classification (a) accuracy and (b) Expected Calibration Error averaged over six folds when training on 10% of the data, comparing T3S (Ours) and U-TAE.
-Figure 10: In‐season classification performance at the 75th‐ and 50th‐percentile cutoffs, corresponding to the end of September and the end of June, respectively, showing (a) accuracy and (b) Expected Calibration Error, averaged over six folds, for T3S (Ours) versus U-TAE.
-Figure 11: Qualitative comparison for the proposed T3S and U-TAE.
-In performance maps, red indicates a wrong prediction while green represents a correct prediction.
-High uncertainty is indicated by yellow, while low is indicated with dark blue.
-6 Discussion
-Our experiments demonstrate that embedding GDD-informed sampling into a deep learning pipeline not only improves crop classification accuracy but also yields better-calibrated uncertainty estimates compared to leading baselines.
-This advantage grows when training data are scarce, highlighting the value of domain expertise under low-data regimes.
-Injecting domain knowledge into a neural network, i.e., inductive bias, constrains the hypothesis space to solutions that ’make sense’ under known physical or biological rules and this prevents the model from assigning high confidence to predictions when evidence is weak or contradictory (Mitchell, (1980), Ahmadi Daryakenari et al., (2024)).
-In our case inductive bias is introduced via GDD-informed sampling.
-By incorporating thermal time alignment, we align satellite time series with actual crop phenology.
-This inductive bias acts as a regularizing prior, steering the model toward meaningful growth stages and away from noisy or cloud-contaminated observations.
-As a result, high-confidence predictions coincide with clear phenological cues, while lower-confidence outputs correctly signal uncertainty, demonstrating reliable probability calibration.
-By combining deep neural networks with agronomic expertise, our approach not only enhances classification accuracy but also empowers the model to recognize and flag its own uncertainty, an essential feature in agricultural decision‐making, where misplaced confidence can have costly consequences.
+Figure 7: In‐season classification performance at 75th‐ and 50th‐percentile cutoffs, corresponding to end of September and June, respectively, showing (a) accuracy and (b) expected calibration error (ECE), averaged over six folds, for T3S (Ours) and U-TAE.
+Low Data Regime.
+In many agricultural settings, annotated crop labels are scarce or expensive to obtain, which makes low-data performance a critical requirement for operational viability.
+To assess robustness under label scarcity, we conducted a six-fold experiment using only 10 % of the training data, randomly sampled once and held constant across methods, and evaluated both T3S and U-TAE on the full test set.
+As shown in Figure 8, both models incur drops in overall accuracy and increases in ECE when labels are limited, but U-TAE’s performance deteriorates far more rapidly.
+Remarkably, T3S trained on just 10 % of the data still outperforms U-TAE trained on the full dataset, and its calibration remains substantially more reliable.
+Discussion.
+The strength of T3S lies in its induced inductive bias: By embedding temperature-informed sampling (the thermal time prior) into the model, we inject domain knowledge that constrains its hypothesis space to solutions that ’make sense’ under known physical or biological rules and prevents overconfident predictions when evidence is weak or conflicting [30, 1].
+In practice, the thermal-time prior helps the model concentrate on critical growth stages and discard noise, such as cloud-contaminated observations, steering it toward meaningful phenological signals, shaping more informative learned representations; thus, boosting performance and producing more reliable confidence estimates, particularly in low data regimes, where inductive biases play the greatest role [57, 25].
+Consequently, the T3S advantage over the baseline is more pronounced under scarce supervision.
+It produces well-calibrated uncertainty estimates: clear phenological cues yield high-confidence predictions, while ambiguous or conflicting inputs appropriately result in low confidence.
+Figure 8: Low data regime performance showing (a) accuracy and (b) expected calibration error (ECE) averaged over six folds when training on 10% of the data, comparing T3S (Ours) and U-TAE.
+Figure 9: Qualitative comparison between T3S and U-TAE.
+Green pixels denote correct predictions, red pixels denote errors, and dark blue indicates low uncertainty, while yellow indicates high uncertainty.
+For best visibility, view the figure zoomed on-screen.
 6 Conclusion
-In this paper, we introduced Thermal Time‐based Temporal
-Sampling (T3S), a simple, model-agnostic method that uses cumulative growing degree days to select relevant observations within satellite time series for crop growth-related applications.
-Based on basic ecophysiological principles of crop growth, it reduces the computational load by removing redundant information and allows to prioritize relevant and undisturbed imagery.
-Although we evaluated T3S with attention-based deep learning architecture, it can be seamlessly integrated with any machine learning model.
-When compared to other state-of-the-art approaches, T3S not only enhances predictive accuracy but also markedly improves uncertainty estimation, an aspect often overlooked in the literature.
-Furthermore, we made evaluation on SwissCrop, a novel, large-scale, multi-year dataset spanning three growing seasons (2021–2023) across Switzerland’s agricultural landscapes, thereby providing a real-world benchmark for compared methods.
-By releasing both SwissCrop and our T3S implementation to the public, we establish a robust, reproducible benchmark for the community, facilitating fair comparisons and driving further innovation in crop‐growth modelling.
-For future work, incorporating additional environmental drivers such as precipitation could further strengthen temporal alignment and robustness across diverse agro-climatic regions.
-8 Declaration of Competing Interest
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-9 Acknowledgments
-Mehmet Ozgur Turkoglu acknowledges support from Agroscope through research funding for the CropmAIpper project.
-Sélène Ledain acknowledges funding from the Swiss Federal Office of Agriculture and Agroscope within the Monitoring of the Swiss Agri-Environmental System (MAUS) program.
-This project was conducted in the context of the Swiss Agricultural Landscape Intelligence platform, funded through an Agroscope Research grant awarded to Helge Aasen.
+We present Thermal Time–based Temporal Sampling (T3S), a model-agnostic approach that leverages cumulative growing degree days to subsample satellite time series in biologically meaningful ways.
+By injecting this thermal-time inductive bias, T3S steers the attention toward biologically relevant periods, reducing redundancy and noise and preventing overconfident predictions under weak or contradictory evidence.
+We evaluated T3S on SwissCrop, a large, multi-year (2021–2023) dataset covering all of Switzerland, and release both the data and our implementation as a public benchmark.
+Compared to state-of-the-art methods, T3S improves both classification accuracy across years and uncertainty estimation, an aspect often overlooked in the literature.
+Since temperature is a main driver for many biological processes, we anticipate that T3S also has advantages for many other applications even beyond plant science.
+And since T3S is model-agnostic, it is applicable to any learning model.
+We leave these explorations for future work.
+Appendix
+Appendix A Training Details
+The model was trained for 100 epochs using a batch size of 8 and a fixed temporal length ($T$) of 24.
+The input image size ($H$ and $W$) is set to $128\times 128$ pixels.
+We employed the Adam optimizer with an initial learning rate of 0.001, scheduled by a OneCycleLR policy with a maximum learning rate of 0.01 and a warm-up phase covering 30% of the total training steps.
+The U-TAE architecture is configured with an encoder having channel widths $[128,128,128,256]$, a symmetric decoder with widths $[32,64,128,256]$ and strided convolutions with kernel size $k=4$, stride $s=2$, and padding $p=1$.
+The temporal attention module employs $n_{\text{head}}=16$ heads and a model dimension of $d_{\text{model}}=256$.
+Input features are standardized to zero mean and unit variance on a per-channel basis.
+We implemented the model in PyTorch and trained it on a single NVIDIA RTX 4090 GPU over roughly three days.
+The full codebase and pretrained checkpoints will be publicly available on GitHub.
+Appendix B Dataset
+In this work, we work with SwissCrop 2021, 2022, 2023, a three-year crop dataset covering the entirety of Switzerland (see Figure 10) from 2021 to 2023.
+SwissCrop combines Sentinel-2 Level-2A bottom-of-atmosphere multi-spectral imagery, averaging 110 timestamps per season, with annual crop‐type labels for each field.
+The dataset includes 50 distinct crop types, each representing the primary crop grown in a field during a given season.
+These labels where provided by farmers to the cantons of Switzerland according to a data model provided by the Federal Office for Agriculture.
+[^1] https://www.blw.admin.ch/de/landwirtschaftliche-kulturflaechen Unlike many existing datasets, e.g.
+Pastis dataset [14], SwissCrop faithfully reflects the pronounced class imbalance found in real‐world farming: a long‐tail distribution in which a handful of major crops dominate while the majority occur only sporadically (see Figure 11).
+[…]
+To complement these remote sensing observations, we incorporate high-resolution gridded climate data from MeteoSwiss, which supplies near-surface temperature fields (minimum, maximum, and average) on 1 km, 2 km, and 5 km grids [11].
+In our study, we use the 1 km daily minimum (TminD) and maximum (TmaxD) temperature series to compute growing degree days (GDD; refer to Section 3.3).
+We make the full SwissCrop dataset publicly available to the research community.
+We hope this resource will facilitate further advances in large-scale, cross-year crop classification, phenology modeling, and earth observation research.
+Figure 10: Region of interest (top left), the SwissCrop dataset (bottom left) and its coverage by data cubes (center), with an example of a single analysis-ready cube (right).
+Figure 11: SwissCrop dataset label distribution for the year 2021. Note the logarithmic scale of the y-axis.
+Appendix C Extended Related Work
+Deep Learning for Crop Mapping from Satellite Imagery.
+Deep learning has dramatically improved crop‐type classification from satellite time series, outperforming traditional methods based on handcrafted features [50].
+Early sequence‐modeling approaches used recurrent architectures to capture phenological dynamics.
+Long short‐term memory (LSTM) networks demonstrated strong performance on multi‐temporal inputs [38], and the convolutional‐RNN (convRNN) version further boosted both accuracy and efficiency [51].
+Attention mechanisms have emerged as a powerful alternative for modeling long‐range temporal dependencies.
+[41] and [15] were the first to apply self‐attention to crop classification from Sentinel-2 data and showed improved results.
+[44] then proposed the lightweight Temporal Attention Encoder (L-TAE), and [14] embedded L-TAE within a U-Net backbone to adaptively weigh temporal features across space and time.
+More recently, fully attention-based Vision Transformer (ViT) adaptations have been explored for crop mapping [2], and [33] leveraged the recently proposed Mamba state‐space architecture to improve spatio‐temporal classification performance.
+To incorporate domain knowledge in deep learning architectures, several works introduce structured priors and hierarchical labels.
+[55] exploited multi‐scale crop class hierarchies within a parameter‐efficient convRNN (i.e, convSTAR [52]) framework, and prototype‐based supervision has been used to integrate class relationships into metric learning [45].
+Positional encodings tailored to agricultural cycles have also been proposed: [31] introduced thermal-time encoding, using cumulative growing degree days, to normalize phenological shifts across regions.
+This is the most relevant research for this paper.
+Moreover, practical deployment demands handling early-season data and irregular revisit intervals.
+[43] added an early classification reward loss to LSTM training, improving accuracy in the season when only partial time series are available.
+[29] leveraged neural ordinary differential equations (NODE) to process noisy and irregularly sampled time series, and [7]’s collect–update–distribute design further addresses sampling irregularity.
+Multi-Year Crop Mapping from Satellite Imagery.
+Multi-year crop mapping faces challenges due to temporal domain shifts such as interannual variability in weather, phenological differences, and sensor calibration issues.
+These factors often degrade the performance of models trained on single-season data when applied across multiple years.
+[54] extended the experiments of [51] to a multi‐year setting, training and testing on data from two different years, and reporting preliminary results.
+[5] employed a Random Forest classifier using dense Sentinel-2 and Landsat-8 time series, monthly Sentinel-1 composites, and environmental variables to evaluate the relative contributions of optical, radar, and environmental data across three meteorologically distinct years (2017–2019).
+[9] demonstrated that reconstructing multi-year Sentinel-2 NDVI time series with Fourier-based fitting and applying Gradient Boosting enables improved crop mapping across different years.
+[13] investigated ways to optimize the use of multi-year samples in a within-season crop classification model.
+[32] presented a Google Earth Engine workflow that fuses field-sample data with multi-year Sentinel-2 time series to map cropping systems on smallholder farms, achieving consistently robust results across diverse landscapes.
+[26] benchmarked deep learning segmentation models (e.g., U-Net and SegFormer [60]) for crop mapping across multiple years, while [47] evaluated convLSTM and convSTAR [52] networks in a similar multi-year context.
+Very recently, [22] shows improved results for cross-year crop mapping combining a U-Net with photometric augmentation, Tversky-Focal loss, and Monte Carlo (MC) Dropout.
+Uncertainty Modeling in Deep Learning.
+Predictive uncertainty is commonly split into aleatoric uncertainty, which stems from inherent data variability such as class overlap or sensor noise, and epistemic uncertainty, which reflects model weight uncertainty due to insufficient coverage of the input space [21].
+Exact Bayesian inference in deep networks is generally intractable, leading to approximate methods like variational inference (e.g., Bayes by Backprop [6] and Markov Chain Monte Carlo techniques such as Stochastic Gradient Langevin Dynamics [59]).
+However, these approaches often struggle to scale to large architectures [18].
+Probabilistic ensembles, particularly deep ensembles, currently represent the gold standard for epistemic uncertainty estimation in deep networks, providing well‐calibrated and robust predictions by aggregating outputs from multiple independently trained models [23].
+However, the need to train and store several large networks incurs substantial computational and memory overhead.
+To address this, implicit ensemble techniques, such as Monte Carlo (MC) Dropout [12], Snapshot Ensembles [20], Masksembles [10], and FiLM-Ensemble [53], approximate ensemble diversity via stochastic weight perturbations or dynamic weight combinations, significantly reducing resource requirements.
+Despite these gains in efficiency, implicit methods generally underperform full ensembles in both predictive accuracy and uncertainty calibration.
+More recently, LoRA-Ensemble has been introduced as a parameter‐efficient alternative for transformer‐based architectures, achieving uncertainty estimates competitive with deep ensembles at a fraction of the cost [19].
+Uncertainty Modeling in Satellite Imagery and Crop Mapping
+Uncertainty quantification has become a standard component of satellite-based land-cover and crop-mapping workflows, as it provides essential information for risk-aware decision-making in agricultural management and ecosystem monitoring.
+Generating per-pixel or per-field uncertainty maps alongside deterministic predictions enables practitioners to identify low-confidence areas and tailor downstream analyses accordingly.
+For example, [24] employs a deep ensemble of convolutional neural networks to estimate global canopy height from sparse LiDAR samples and multi-sensor satellite imagery, producing uncertainty intervals that reflect both model and data variability.
+[4] extends this ensemble approach to country-scale forest structure estimation at 10m resolution, demonstrating that dense uncertainty fields help flag regions where auxiliary field data are most needed.
+In the context of crop yield and type mapping, several recent studies have adopted stochastic inference techniques.
+[49] introduces an attention-guided multi-level crop network for winter wheat yield estimation at the county scale, applying MC Dropout during inference to derive spatially explicit uncertainty maps of yield forecasts.
+Similarly, [22] leverages MC Dropout in a U-Net segmentation framework to quantify per-pixel uncertainty in annual cropland masks across multiple years in Ghana, using the resulting confidence estimates to discard unreliable predictions.
+[27] integrates MC Dropout in a ViT–ResNet feature encoder to produce field-level uncertainty scores, filtering out low-confidence labels to improve overall mapping precision.
+Beyond dropout-based methods, [3] proposes a probabilistic crop-type mapping method in which an ensemble of classifiers, each trained with perturbed inputs and hyperparameters on Sentinel-1, Sentinel-2, and Landsat-8 data, generates multiple crop maps over Germany; per-pixel frequency counts across the ensemble yield robust probability distributions and uncertainty metrics.
+[28] refines the 30m cropland data by harnessing its built-in Random Forest confidence layer: low-certainty regions are iteratively reclassified using targeted ancillary datasets, resulting in a more accurate and self-aware mapping product.
+Collectively, these approaches demonstrate the critical role of uncertainty modeling in improving the reliability and interpretability of satellite-driven crop and canopy mapping systems.